### v1

# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

Abstract
In this paper, we study the Lagrangian aspects of the inhomogeneous incompressible Euler equation (IIE in short). We establish a geodesic description of this equation and discuss the associated geometric structures. We also find the derivation of IIE from the Hamilton-Pontryagin action principle and derive the corresponding Lagrangian formulation. Appealing to this Lagrangian perspective, we prove Lagrangian analyticity of IIE.
1 Introduction
The geometry of hydrodynamics has been extensively studied since the seminal work of Arnold [Ar66] and Ebin-Marsden [EM70], where the motion of perfect fluids in a smooth bounded domain $D$ is interpreted as geodesic in the infinite dimensional Riemannian manifold $\mathsf{SDiff}(D)$ of volume-preserving diffeomorphisms of $D$, which is equipped with the right-invariant Riemannian metric given by the kinetic energy of the fluid. This geometric perspective has been extensively generalized since then; we refer to [AK98][HMR98][MR99] for further development of this topic, and [KMM20] for a recent survey on geometric hydrodynamics.
From a mechanical viewpoint, Arnold’s geodesic formulation is a natural infinite-dimensional generalization of the classical least action principle in Lagrangian mechanics. Flipping to the Hamiltonian side, we have an analogous Hamilton-Pontryagin type variational principle for hydrodynamic models (see [BCHM00][Ho21]), which leads to the Lagrangian-Eulerian formulation of fluid equations [Co01a]. We remark that the above perspective admits a natural generalization which effectively applies to viscous hydrodynamic models. We refer to [CI08] for the Navier-Stokes case and to [MP25] for an extended discussion. However, all the above discussions are restricted to fluid models with mechanical Lagrangians, where kinetic and potential energy are separated.
A natural example which lies out of the zoo of mechanical Lagrangian is the following inhomogeneous version of the incompressible Euler equation (IIE in short), which is an extensively studied model in past decades:
$\left\{\begin{aligned} &\partial_{t}(\rho u)+\nabla\cdot(\rho u\otimes u)+\nabla p=0\\ &\nabla\cdot u=0\\ &\partial_{t}\rho+u\cdot\nabla\rho=0\end{aligned}\right$. (1)
The above equation describes fluids with variable density, transported by its own incompressible velocity field. The study of the above equation was initiated in the celebrated monograph [[Lio96]], see also [Da06][[Da10]][[DF11]]. The main focus of the works listed above is to address well-posedness issues in various functional settings, analyzing IIE from an Eulerian viewpoint using delicate Harmonic analysis tools. Contrary to the well-known Yudovich theorem of 22-$d$ Euler’s equation, the global well-posedness issue of IIE in 22d is still open. There are also works devoted to studying the Beale-Kato-Majda (cf.[[BKM84]] )type blow up/continuation criteria for IIE in the past few years, we refer the readers to [[BLS20]][[Fa25]].
However, different from the analytical side, the geometric structure and Lagrangian perspective of the inhomogeneous Euler equation is much less studied. To the author’s knowledge, the only relevant literature in this direction is [[Mar76]], which addressed the geodesic formulation of IIE on $\mathsf{SDiff}(D)$, as a generalization of [EM70]. To fill the gap in the literature, in this paper we establish a Lagrangian approach to the analysis of IIE, which stems from its geometric structure. We discuss the Hamiltonian formulation of the equation and derive a Lagrangian-Eulerian representation of the solution of IIE. The advantages of such Lagrangian perspective are two-folds: First, the Lagrangian formulation motivates us to find a vorticity formulation of IIE, which totally eliminates the pressure; Second, benefited from the Lagrangian perspective, we obtain an explicit recursive formula for Taylor coefficients (in time) of Lagrangian displacement, and accordingly, demonstrate Lagrangian analyticity of the solution.
The structure of this paper is organized as follows: In section 2, we study the geodesic formulation of IIE with respect to a weighted Riemannian metric on $\mathsf{SDiff}(D)$, where we interpret the weighted Leray projector as the second-fundamental form. The sectional curvature and Lagrangian stability are also discussed. In section 3, we derive IIE from a Hamilton-Pontryagin type variational principle, and obtain a Lagrangian-Eulerian formulation of IIE as a Lagrangian formulae for solution of the Hamiltonian ODE given by the variational principle. We also derive the vorticity formulation from our geometric construction, and recover the local existence results in our framework. Finally in section 4, using our Lagrangian representation, we prove the analyticity of Lagrangian trajectories of the IIE.
In this paper, we will use the letter $C$ to denote generic positive constants which may change from line to line. $A\lesssim B$ means there exists $C>0$ such that $A\leq CB$.
Acknowledgements
The author thanks T.Drivas for pointing out the ongoing work [DG26+] and for interesting discussions. The author thanks A.Mazzucato and A.Novikov for stimulating discussions and valuable suggestions. A. Pan was partially supported by the US National Science Foundation grants DMS-1909103 and DMS-2206453 (PI Mazzucato).
2 Geometry of $\mathsf{SDiff}(D)$ with Weighted Metric
In this section, we study the Riemannian structure of $\mathsf{SDiff}(D)$ which naturally fits the inhomogeneous incompressible Euler equation. Our presentation in this section mainly follows and generalizes the discussions in the unpublished note [DrWS25], which generalizes constrained geodesic, second fundamental form and sectional curvature of level set submanifold from finite to infinite dimension (The $\mathsf{SDiff}(D)$ setting).
We start with the well-known Jacobi formula which will be repeatedly used in this section.
Lemma 2.1
Let $A:[0,1]\to\mathsf{GL}(d)$ be a smooth curve in the general linear group $\mathsf{GL}(d)$, then we have:
$\frac{d}{dt}\det(A_{t})=\mathsf{Tr}(A^{-1}\dot{A})\det(A)$ (2)
$\frac{d}{dt}A_{t}^{-1}=-A_{t}^{-1}\dot{A}_{t}A_{t}^{-1}$ (3)
Proof. To show (2), let $\{M_{ij}\}$ be minors of $A=\{a_{ij}\}$, we expand for each fixed $i$:
$\det(A)=\sum_{j=1}^{d}(-1)^{i+j}M_{ij}a_{ij}$
Using the fact that $M_{ij}$ is independent of $a_{ij}$, we have
$\frac{\partial\det(A)}{\partial a_{ij}}=(-1)^{i+j}M_{ij},\quad\frac{d}{dt}\det(A)=\sum_{i,j=1}^{d}\frac{\partial\det(A)}{\partial a_{ij}}\frac{da_{ij}}{dt}=\sum_{i,j=1}^{d}(-1)^{i+j}M_{ij}\frac{da_{ij}}{dt}$
Since we have
$(A^{-1})_{ij}=\frac{1}{\det(A)}(-1)^{i+j}M_{ij}$
We conclude
$\frac{d}{dt}\det(A)=\det(A)\sum_{i,j=1}^{d}(A^{-1})_{ij}\frac{da_{ij}}{dt}=\mathsf{Tr}(A^{-1}\dot{A})\det(A)$
(3) is a simple consequence of the chain rule:
$0=\frac{d}{dt}(A_{t}A_{t}^{-1})=\dot{A}_{t}A_{t}^{-1}+A_{t}\frac{d}{dt}A_{t}^{-1}$
Rearranging the above, we obtain (3). ∎
2.1 The Riemannian Structure
Let $D$ be either $\mathbb{R}^{d}$ or $\mathbb{T}^{d}$, which is the physical domain of our fluids, with natural decaying/periodic boundary conditions. Throughout this paper, we denote by $\mathsf{Diff}(D)$($\mathsf{SDiff}(D)$) the Lie group of (volume-preserving) diffeomorphisms in $D$. We denote by $\mathfrak{X}(D)$ ($\mathfrak{X}_{\sigma}$) the space of (divergence-free) smooth vector fields on $D$. We remark that the formulation below works in the setting of $H^{s+1}$ Sobolev regularity, with $s>d/2$. Here, we only consider the smooth case for simplicity.
Our geometric description relies on the interpretation of $\mathsf{SDiff}(D)$ as a level set submanifold of the full diffeomorphism group $\mathsf{Diff}(D)$ with infinite dimension and co-dimension, and the inhomogeneous Euler equation will be derived as geodesic on the submanifold, with respect to some weighted metric. This formalism closely follows [DrWS25]. (See also the ongoing work [DG26+], dealing with dynamics on rapidly oscillating constraint submanifold in infinite dimensions.)
More precisely, we reformulate
$\mathsf{SDiff}(D)=\bigcap_{a\in D}\{X\in\mathsf{Diff}(D):f_{a}(X)=0\},\quad f_{a}(X)=\langle\det(\nabla X)-1,\delta_{a}\rangle_{*}.$
Here, the duality pairing $\langle\cdot,\cdot\rangle_{*}$ is the Riesz pairing between $C_{0}$ functions and measures.
Now for $X\in\mathsf{SDiff}(D)$, to see the tangential direction of the level sets of $f_{a}$, we apply (2) to compute
$\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}f_{a}(X+\varepsilon\delta X)=\langle\mathsf{Tr}((\nabla X)^{-1}\nabla\delta X)\det(\nabla X),\delta_{a}\rangle_{*}$
$\nabla\cdot(\delta X\circ X^{-1})](X(a))=0.$
Which implies that
$T_{X}\mathsf{SDiff}(D)=\{v\circ X:\nabla\cdot v=0\}.$
Now we move on to define the Riemannian metric on $\mathsf{SDiff}(D)$. We denote by $\mathsf{Dens}(D)$ the space of smooth densities in $D$. Fix a reference density $\rho_{0}\in\mathsf{Dens}(D)$ which is required to avoid concentration or vacuum, that is, satisfying the following assumption in this paper:
Assumption (A):
There exists $C_{1}>1$ such that
$0<C_{1}^{-1}\leq\rho_{0}(a)\leq C_{1}<\infty\quad\text{ for all }a\in D.$
Define the following weighted metric:
$\mathsf{G}_{X}(\xi,\eta)=\int_{D}(\xi,\eta)(x)\rho_{0}(x)dx=\int_{D}(\xi\circ X^{-1},\eta\circ X^{-1})\rho(x)dx.$ (4)
here $\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A$. Clearly $\rho$ is a rearrangement of $\rho_{0}$, so the assumption (A) is preserved under action of $\mathsf{SDiff}(D)$ on $\mathsf{Dens}(D)$.
Under the above metric, the normal space $N_{X}\mathsf{SDiff}(D)=T_{X}^{\perp}\mathsf{SDiff}(D)$ is given by
$N_{X}\mathsf{SDiff}(D)=\{\rho_{0}^{-1}\nabla\varphi(X):\varphi\in H^{1}(D)\}.$ (5)
Hence, we can define the following map that generates normal vectors:
$\mathcal{N}_{X}:\varphi\in H^{1}(D)\to\rho_{0}^{-1}\nabla\varphi(X)=\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(\varphi\circ X)\in N_{X}\mathsf{SDiff}(D).$ (6)
We now extend the domain of the above normal generator $\mathcal{N}$ to distributions. Let $\mathcal{M}$ be the set of Borel measures with finite variation, for any $\mu\in\mathcal{M}$ we define the following co-normal functional
$\mathcal{N}_{X}^{*}(\mu):\delta X\in T_{X}\mathsf{Diff}(D)\to\langle\mu,\nabla\cdot(\delta X\circ A)\rangle_{*}.$ (7)
which satisfies the annihilating property $T_{X}\mathsf{SDiff}(D)\subset\mathsf{Ker}(\mathcal{N}_{\mu}^{X})$.
Moreover, in this context, the musical isomorphism (inertia operator in the Language of mechanics) is $\alpha\to\rho_{0}^{-1}\alpha$, since for any $1$-form $\alpha$ we have
$\langle\delta X,\alpha\rangle_{*}=\mathsf{G}_{X}(\alpha_{\sharp},\delta X)=\langle\delta X,\alpha_{\sharp}\rangle_{L_{\rho_{0}}^{2}}.$
Which implies that $\alpha_{\sharp}=\rho^{-1}\alpha$. Hence, the vector representative $\mathcal{N}_{X}$ of covector $\mathcal{N}_{X}^{*}$ is:
$\mathcal{N}_{X}(\mu):=\rho_{0}^{-1}\mathcal{N}_{X}^{*}(\mu)=-\rho_{0}^{-1}\nabla\mu(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(X_{*}\mu).$
which should be interpreted in the distributional sense.
The following proposition is a consequence of standard computation:
Proposition 2.2
The following statements hold:
  • (i) For any $a\in D$, the gradient $\mathsf{grad}f_{a}$ of $f_{a}$ is:
    $\mathsf{grad}f_{a}(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla\delta_{a}=\mathcal{N}_{X}(\delta_{X(a)})=\mathcal{N}_{X}(\delta_{a}\circ A)$ (8)
  • (ii) For any $a\in D$, the Hessian $\mathsf{Hess}f_{a}$ of $f_{a}$ is
    $\mathsf{Hess}f_{a}(X)(\xi,\eta)=-\mathsf{Tr}[\nabla(\xi\circ A)\nabla(\eta\circ A)](X(a)),\quad\forall\xi,\eta\in T_{X}\mathsf{SDiff}(D)$ (9)
  • (iii) For any $a,a^{\prime}\in D$, the Gram operator $(Jf)_{a,a^{\prime}}$ is
    $(Jf)_{a,a^{\prime}}:=\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\mathsf{grad}f_{a^{\prime}}(X))=\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(a)).$ (10)
    While the inverse $(Jf)_{a,a^{\prime}}^{-1}$ reads
    $(Jf)_{a,a^{\prime}}^{-1}(X)=G_{\rho}(X(a),X(a^{\prime})).$ (11)
    where $\Delta_{\rho}=\nabla\cdot(\rho^{-1}\nabla)$ and $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$.
Proof. The proof follows from straightforward computation. For (I) we have
$\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\delta X)=\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}f_{a}(X_{\varepsilon})=\langle\nabla\cdot(\delta X\circ A)\circ X,\delta_{a}\rangle_{*}.$
$\langle\delta X\circ A,\nabla(\delta_{a}\circ A)\rangle_{*}=-\langle\delta X,(\nabla^{*}X)^{-1}\nabla\delta_{a}\rangle_{*}$
So we have for any $\delta X\in T_{X}\mathsf{SDiff}(D)$:
$\int_{D}(\mathsf{grad}f_{a}(X),\delta X)\rho_{0}(x)dx=-\int_{D}\delta X\cdot(\nabla^{*}X)^{-1}\nabla\delta_{a}dx.$
Hence, we conclude
$\mathsf{grad}f_{a}(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla\delta_{a}$ (12)
which completes the proof of (I). For (II), picking any vector fields $u,v\in\mathfrak{X}_{\sigma}$ and denoting $X_{\varepsilon}:=X+\varepsilon v\circ X$, we have:
$\mathsf{Hess}f_{a}(X)(u\circ X,v\circ X)=\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}\mathsf{G}_{X}(\mathsf{grad}f_{a}(X_{\varepsilon}),u\circ X).$
$\int_{D}-\rho_{0}^{-1}\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}(\nabla^{*}X_{\varepsilon})^{-1}\nabla\delta_{a}\cdot u(X)\rho_{0}(x)dx$
$\int_{D}(\nabla^{*}X)^{-1}\nabla^{*}(v\circ X)(\nabla^{*}X)^{-1}\nabla\delta_{a}\cdot u(X)dx.$
$\int_{D}\nabla\delta_{a}(\nabla Au\cdot\nabla v)(X)dx$
$\int_{D}\delta_{a}\nabla\cdot[A_{\sharp}(u\cdot\nabla v)]dx$
$\nabla\cdot(u\cdot\nabla v)X(a)=-\mathsf{Tr}(\nabla u\cdot\nabla v)(X(a)).$
which yields (II). Now for (III), we compute:
$\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\mathsf{grad}f_{a^{\prime}}(X))=\int_{D}(\nabla^{*}X)^{-1}\nabla\delta_{a}(x)(\nabla^{*}X)^{-1}\nabla\delta_{a^{\prime}}(x)\rho_{0}^{-1}(x)dx$
$\int_{D}\nabla(\delta_{a}\circ A)(X)\nabla(\delta_{a^{\prime}}\circ A)(X)\rho_{0}^{-1}(x)dx$
$\int_{D}\nabla(\delta_{a}\circ A)\nabla(\delta_{a^{\prime}}\circ A)\rho_{0}^{-1}\circ Adx$
$\int_{D}(\delta_{a}\circ A)\nabla\cdot[\rho^{-1}\nabla(\delta_{a^{\prime}}\circ A)]$
$\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(a)).$
Meanwhile, the inverse $(Jf)_{a,a^{\prime}}^{-1}$ satisfies:
$\int_{D}(Jf)_{a,z}^{-1}(X)(Jf)_{z,a^{\prime}}(X)dz=\delta(a-a^{\prime}),$
which implies
$\delta(a-a^{\prime})=-\int_{D}(Jf)_{a,z}^{-1}(X)\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(z))dz$
$\Delta_{\rho}[(Jf)_{a,A(z)}^{-1}(X)](X(a^{\prime})).$
Hence, we conclude:
$\Delta_{\rho}[(Jf)_{a,A(z)}^{-1}(X)]=\delta(X(a)-z).$
As a consequence, we have
$(Jf)_{a,a^{\prime}}^{-1}(X)=-G_{\rho}(X(a),X(a^{\prime})),$
where $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$.
2.2 Geodesic Equation
Now let $X$ be a geodesic on infinite dimensional Riemannian manifold $(\mathsf{SDiff}(D),\mathsf{G})$. With the above computation, D’Alembert principle implies that $\ddot{X}\in T_{X}^{\perp}\mathsf{SDiff}(D)$, hence there exists a scalar function $\lambda:[0,T]\times D\to\mathbb{R}$ such that for any $x\in D$:
$\rho_{0}(x)\ddot{X}(x)=\int_{D}\lambda(t,z)(\nabla^{*}X)^{-1}(x)\nabla_{z}\delta_{x}(z)dz$ (13)
$(\nabla^{*}X)^{-1}(x)\int_{D}\underbrace{\nabla_{z}(\delta_{z}(x)\lambda(t,z))}_{=0\quad a.e.}-\delta_{z}(x)\nabla\lambda(t,z)dz$
$(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(z)\nabla(\underbrace{\lambda\circ A}_{:=p})(X(z))\delta_{z}(x)dz$
$(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(z)\nabla p(X(z))\delta_{x}(z)dz=-\nabla p(t,X).$
Which, in Eulerian coordinates, reads (1).
We take a closer look at the Lagrange multiplier $\lambda$. Since for any $a\in D$ we have:
$\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\dot{X})=0.$ (14)
Differentiating (14) in time and denoting by $u=\dot{X}_{t}\circ X_{t}^{-1}$, we have
$0=-\frac{d}{dt}\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\dot{X}_{t}(x)dx.$
$\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\ddot{X}_{t}(x)dx+\int_{D}(\nabla^{*}X_{t})^{-1}\nabla^{*}(\dot{X}_{t}(x))(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}\cdot\dot{X}_{t}(x)dx$
$\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\ddot{X}_{t}(x)dx+\int_{D}\nabla\delta_{a}(x)(A_{t})_{\sharp}(u\cdot\nabla u)dx$
$\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\ddot{X}_{t}(x)dx-\int_{D}\delta_{a}(x)[\nabla\cdot(u\cdot\nabla u)](X_{t}(x))dx.$
Plugging (13) into the above equality, we have the following.
$\nabla\cdot(u\cdot\nabla u)](X_{t}(a))=-\int_{D}(\nabla^{*}X_{t})^{-1}(x)\nabla\delta_{a}(x)\int_{D}\rho_{0}^{-1}(x)\lambda(z)(\nabla^{*}X_{t})^{-1}(x)\nabla_{z}\delta_{x}(z)dzdx$
$\int_{D}\int_{D}\nabla(\delta_{a}\circ A_{t})(X_{t}(x))\rho_{0}^{-1}(x)\lambda(z)(\nabla^{*}X_{t})^{-1}(x)\nabla_{z}\delta_{x}(z)dzdx$
$\int_{D}\int_{D}\nabla(\delta_{a}\circ A_{t})\rho_{t}^{-1}(x)\nabla^{*}A_{t}(x)\nabla\lambda(z)\delta_{A_{t}(x)}(z)dzdx$
$\int_{D}\nabla(\delta_{a}\circ A_{t})\rho_{t}^{-1}(x)\nabla(\lambda\circ A_{t})(x)dx$
$\nabla\cdot(\rho_{t}^{-1}\nabla p)(X_{t}(a)),\quad p:=\lambda\circ A_{t}.$
Hence, we conclude:
$p=-\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u),\quad\lambda=p\circ X.$ (15)
The above geodesic equation in Eulerian coordinate reads:
$\rho(\partial_{t}u+u\cdot\nabla u)-\nabla\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u)=0$
Together with $\nabla\cdot u=0$ and $\rho$ transported by $u$, we readily verify that IIE (1) is the geodesic equation on $\mathsf{SDiff}(D)$ with our weighted metric $\mathsf{G}$.
2.3 Second Fundamental Form and Curvature Tensor
Viewing $\mathsf{SDiff}(D)$ as a submanifold of $\mathsf{Diff}(D)$ of infinite dimension and codimension, it makes sense to talk about the second fundamental form of $\mathsf{SDiff}(D)$, which admits a concrete description as the projection of the convective acceleration onto the normal direction. To see this, we define for any given $\rho\in\mathsf{Dens}(D)$ the following operator
$\mathbf{Q}_{\rho}:w\to\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot w$ (16)
The next lemma shows that $\mathbf{Q}_{\rho}$ is an orthogonal projector in weighted $L^{2}$-sense.
Proposition 2.3
Let $\rho\in\mathsf{Dens}(D)$ and $w\in\mathfrak{X}(D)$, then the following statements hold:
  • (i) There exists a unique $v\in\mathfrak{X}_{\sigma}$ and $p\in C^{\infty}(D)$ such that
    $w=\rho v+\nabla p$ (17)
    moreover, $v=(\mathbf{Id}-\mathbf{Q}_{\rho})(\rho^{-1}w)=:\mathbf{P}_{\rho}(\rho^{-1}w)$
  • (ii) Let $\mathcal{H}_{\rho}:=\{\rho^{-1}\nabla\varphi:\varphi\in C^{\infty}(D)\}$, then we have $\mathsf{Ker}(\mathbf{Id}-\mathbf{Q}_{\rho})=\mathcal{H}_{\rho}$.
  • (iii) $\mathsf{Ran}(\mathbf{P}_{\rho})\perp\mathsf{Ker}(\mathbf{P}_{\rho})$ in the sense of $L_{\rho}^{2}$.
Proof. (i) Given $w\in\mathfrak{X}(D)$, for any given $\rho\in\mathsf{Dens}(D)$ we write
$w=\rho[\rho^{-1}w-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)]+\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)=\rho\mathbf{P}_{\rho}(\rho^{-1}w)+\rho\mathbf{Q}_{\rho}(\rho^{-1}w)$ (18)
Let $v=\mathbf{P}_{\rho}(\rho^{-1}w)$ and $p=\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)$ we find the decomposition. Moreover, assume $w=\rho v^{\prime}+\nabla p^{\prime}$ for some $v^{\prime}\in\mathfrak{X}(D)$ and $p^{\prime}\in C^{\infty}(D)$, then $p^{\prime}$ can be recovered from solving the following elliptic problem:
$\Delta_{\rho}p^{\prime}=\nabla\cdot(\rho^{-1}\nabla p^{\prime})=\nabla\cdot(\rho^{-1}w)-\underbrace{\nabla\cdot v^{\prime}}_{=0}=\nabla\cdot(\rho^{-1}w).$
Ellipticity of $\Delta_{\rho}$ implies the uniqueness of solution $p^{\prime}$, hence the uniqueness of $v^{\prime}$.
(ii) Picking any $\varphi\in C^{\infty}(D)$, we compute $\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)$ as follows:
$\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)=\rho^{-1}\nabla\varphi-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}\nabla\varphi)=0.$
Which guaranties $\mathcal{H}_{\rho}\subset\mathsf{Ker}(\mathbf{P}_{\rho})$. Conversely, assume we have $\mathbf{P}_{\rho}\xi=0$, then
$0=\nabla\times(\rho\mathbf{P}_{\rho}\xi)=\nabla\times(\rho\xi)-\nabla\times(\nabla\Delta_{\rho}^{-1}\nabla\cdot\xi)=\nabla\times(\rho\xi).$
which ensures that $\xi\in\mathcal{H}_{\rho}$ and completes the proof of (ii).
(iii) The orthogonality simply follows from the fact that $\mathsf{Ran}(\mathbf{P}_{\rho})\subset\mathfrak{X}_{\sigma}(D)$, while for generic $v\in\mathfrak{X}_{\sigma}(D)$ we have:
$\langle v,\rho^{-1}\nabla\varphi\rangle_{L_{\rho}^{2}}=\langle v,\nabla\varphi\rangle_{2}=0.$
The above proposition gives a dynamical decomposition of vector fields into vertical ($\mathfrak{X}_{\sigma}$) and horizontal ($\mathcal{H}_{\rho}$) directions, from the ambient $\mathsf{Diff}(D)$ viewpoint.
We now connect the second fundamental form $\Pi$ of $\mathsf{SDiff}(D)$ w.r.t metric $\mathsf{G}$ and the above weighted projector $\mathbf{Q}_{\rho}$:
Theorem 2.4
Then the second fundamental form $\Pi_{X}$ of $\mathsf{SDiff}(D)$ at $X$ is given by
$\Pi_{X}(u\circ X,v\circ X)=\mathbf{Q}_{\rho}(u\cdot\nabla v),\quad\forall u,v\in\mathfrak{X}_{\sigma}(D)\text{ and }X\in\mathsf{SDiff}(D).$ (19)
Proof. Formula (19) follows from straightforward computation. By definition:
$\Pi_{X}(u\circ X,v\circ X)(x)=\int_{D\times D}(Jf)_{y,z}^{-1}(X)\mathsf{Hess}f_{z}(X)(u\circ X,v\circ X)\mathsf{grad}f_{y}(X(x))dydz$
$\int_{D}[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(y))\rho_{0}^{-1}(x)(\nabla^{*}X(x))^{-1}\nabla_{y}\delta_{x}(y)dy$
$\rho_{0}^{-1}(x)(\nabla^{*}X)^{-1}(x)\int_{D}[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(y))\nabla_{y}\delta_{x}(y)dy$
$\rho_{0}^{-1}(x)(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(y)\nabla[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(y))\delta_{x}(y)dy$
$\rho_{0}^{-1}(x)\nabla[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(x))$
$(\rho^{-1}\nabla[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)])(X(x))=\mathbf{Q}_{\rho}(u\cdot\nabla v)(X).$
Remark 2.5
Above $\Pi_{X}$ is symmetric on $T_{X}\mathsf{SDiff}(D)\times T_{X}\mathsf{SDiff}(D)$, since
$\Pi_{X}(u\circ X,v\circ X)-\Pi_{X}(v\circ X,u\circ X)=\mathbf{Q}_{\rho}([u,v])(X)=0.$
Here we use the fact that $\nabla\cdot[u,v]=0$ provided $u,v\in\mathfrak{X}_{\sigma}(D)$.
With the above explicit formula for $\Pi$, we can now compute the curvature tensor as follows: For tangent vector $\xi=u\circ X,\eta=v\circ X\in T_{X}\mathsf{SDiff}(D)$ with $u,v\in\mathfrak{X}_{\sigma}(D)$, we have
$\mathcal{C}_{\xi,\eta}=\int_{D}\Pi_{X}(\xi,\xi)(z)\cdot\Pi_{X}(\eta,\eta)(z)-\lvert\Pi_{X}(\xi,\eta)(z)\rvert^{2}\rho_{0}(z)dz$
$\int_{D}[\mathbf{Q}_{\rho}(u\cdot\nabla u)\mathbf{Q}_{\rho}(v\cdot\nabla v)-\lvert\mathbf{Q}_{\rho}(u\cdot\nabla v)\rvert^{2}]\rho(z)dz$
In view of the above curvature formula, we recover the Lagrangian instability of shear flows, as steady states of the homogeneous Euler equation.
Remark 2.6
The sign of the above curvature tensor fails to detect Lagrangian instability of more general class of steady states due to the following reason: Pressureless steady states of homogeneous Euler are essentially shear flows: Let $X$ be the flow of pressureless steady state $u$, by balance of momentum we have for any initial position $a\in D$:
$\ddot{X}(t,a)=(\partial_{t}u+u\cdot\nabla u)(X_{t}(a))=0$
which implies that streamlines are all straight lines.
2.4 A Gibbs-Appell Formulation
Different from the celebrated Lagrange’s least action principle/Hamilton’s stationary action principle in analytical mechanics, there is a much lesser popular principle/formalism in classic mechanics, known as Gibbs-Appell formulation, which connects to the so-called Gauss’s principle of least curvature and Hertz’s principle of least constraint (see [[Lew96]]). The natural playground for this formulation is a mechanical system with ideal constraints, which is compatible with the nature of incompressible fluids. We refer to [[TG23]] for the study of the Navier-Stokes equation from this point of view. Instead of minimizing Lagrangian action under the constraint of fixed endpoints, the Gibbs-Appell principle minimizes the so-called Appellian functional, which is a quadratic function in acceleration, at each instant.
In the context of IIE, we prove the following:
Theorem 2.7
Define the Appell function
$\mathscr{A}(\ddot{X})=\frac{1}{2}\int_{D}\lvert\ddot{X}(t,a)\rvert\rho_{0}(da).$ (20)
Then minimizer $X_{*}$ of $\mathscr{A}$ at each instant $t\in[0,T]$, subject to the constraint $X\in\mathsf{SDiff}(D)$, is the solution of the following Newton’s equation:
$\rho_{0}\ddot{X}=-\nabla p(t,X),\quad u=\dot{X}\circ X^{-1},\quad p=-\Delta_{X_{\sharp}\rho_{0}}^{-1}\nabla\cdot(u\cdot\nabla u).$ (21)
Which is the Lagrangian version of (1).
Proof. Thanks to the Lemma 2.1, differentiating constraint $X(t)\in\mathsf{SDiff}(D)$ twice leads to the following second-order condition:
$\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\dot{X}(t,a)]^{2}+\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\ddot{X}(t,a)=0$ (22)
Introducing the Lagrange multiplier $\chi(t,a)$ of the above condition, we define
$U(X):=\sup_{\lambda}\int_{D}\chi(t,a)\big[-\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\dot{X}(t,a)]^{2}+\mathsf{Tr}[(\nabla X)^{-1}(t,a)\nabla\ddot{X}(t,a)]\big]da$
Clearly $U(X)=0$ when $X\in\mathsf{SDiff}(D)$ and $U(X)=\infty$ if $X\notin\mathsf{SDiff}(D)$. Now consider the following augmented Appellian functional:
$\tilde{\mathcal{A}}(\ddot{X}(t),\dot{X}(t),X(t))=\int_{D}\frac{1}{2}\lvert\ddot{X}(t,a)\rvert^{2}\rho_{0}(a)+U(X(t,a))da$ (23)
and associated saddle-point problem:
$\inf_{X\in\mathsf{SDiff}(D)}\int_{D}\frac{1}{2}\lvert\ddot{X}(t,a)\rvert^{2}\rho_{0}(a)da=\inf_{X\in\mathsf{SDiff}(D)}\int_{D}\frac{1}{2}\lvert\ddot{X}(t,a)\rvert^{2}\rho_{0}(a)+U(X(t,a))da.$
$\inf_{X\in\mathsf{SDiff}(D)}\sup_{\chi}\int_{D}\frac{1}{2}\lvert\ddot{X}\rvert^{2}\rho_{0}(a)+\chi\big(-\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}+\mathsf{Tr}[(\nabla X)^{-1}\nabla\ddot{X}]\big)da.$
The above function is quadratic in $\ddot{X}$ and linear in $\chi$, since we notice that the linear term in $\ddot{X}$ can be rewritten as
$\int_{D}\chi(t,a)\mathsf{Tr}[(\nabla X)^{-1}\nabla\ddot{X}](a)da$
$\int_{D}\chi(t,a)\mathsf{Tr}[\nabla(\ddot{X}\circ A)(X(t,a))]da$
$\int_{D}\chi(t,A(a))\nabla\cdot(\ddot{X}\circ A)(a)da$
$\int_{D}\nabla(\chi\circ A)(X(a))\ddot{X}(a)da.$
Thus, the original saddle-point problem can be reformulated as follows:
$\inf_{X\in\mathsf{SDiff}(D)}\sup_{\chi}\int_{D}\frac{1}{2}\lvert\ddot{X}\rvert^{2}\rho_{0}(a)+\nabla(\chi\circ A)(X(a))\ddot{X}(a)+\chi\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}da.$ (24)
Now, since (24) is quadratic in $\ddot{X}$ and linear in $\chi$, we swap the order of inf and sup and consider
$\sup_{\chi}\inf_{X\in\mathsf{SDiff}(D)}\int_{D}\frac{1}{2}\lvert\ddot{X}\rvert^{2}\rho_{0}(a)+\nabla(\chi\circ A)(X(a))\ddot{X}(a)+\chi\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}da.$
Thanks to the quadratic formula, maximizing in $X$ yields
$\ddot{X}(t,a)=-\frac{\nabla(\chi\circ A)(X(a))}{\rho_{0}}$ (25)
Now we aim to find
$\sup_{\chi}\int_{D}-\frac{1}{2\rho}\lvert\nabla(\chi\circ A)\rvert^{2}+(\chi\circ A)\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}(A(a))da$ (26)
Denote by $\tilde{\chi}=\chi\circ A$, simple algebraic manipulation yields
$\int_{D}\tilde{\chi}\mathsf{Tr}[(\nabla X)^{-1}\nabla\dot{X}]^{2}(A(a))da=\int_{D}\mathsf{Tr}[(\nabla u)^{2}(a)]\tilde{\chi}(a)da=\int_{D}\nabla\cdot(u\cdot\nabla u)(a)\tilde{\chi}(a)da.$
and
$\int_{D}-\frac{1}{2\rho}\lvert\nabla(\chi\circ A)\rvert^{2}da=-\frac{1}{2}\int_{D}\rho^{-1}\nabla\tilde{\chi}\cdot\nabla\tilde{\chi}da=\frac{1}{2}\int_{D}\Delta_{\rho}\tilde{\chi}\cdot\tilde{\chi}da.$
Hence, (26) becomes a quadratic problem in $\tilde{\chi}$. The maximizer is computed similarly as follows:
$\Delta_{\rho}\tilde{\chi}=-\nabla\cdot(u\cdot\nabla u)\quad\Longrightarrow\quad\tilde{\chi}=-\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u)=p.$ (27)
Hence, we combine (25) and (27) and conclude the desired result. ∎
3 Hamiltonian and Vorticity Formulation
In this section, we analyze how the above geometric picture flips to the Hamiltonian side. Recall that the Hamiltonian formulation of the motion of ideal fluids reads coadjoint motion of one-form on $\mathfrak{X}_{\sigma}^{*}(D)$, lies in the framework of general Lie-Poisson equations [MR99], while the Leray projector essentially becomes the musical isomorphism from cotangent to tangent vectors. Identifying $\mathfrak{X}_{\sigma}(D)$ with the space of scalar(vector, resp.) potentials by $\nabla^{\perp}(\nabla\times$,resp.) in $2(3$-d,resp.), we obtain the vortex dynamics. Here we extend the above geometric picture to the inhomogeneous setting. A Hamilton-Pontryagin type variational principle and a companion Lagrangian-Eulerian formulation is established. We then establish the vorticity formulation and appreciate such formulation to prove local existence of strong solutions.
3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
We now lift ourselves to the cotangent bundle of $\mathsf{SDiff}(D)$ where the Hamiltonian equation naturally lives and we’ll derive it in our setting. To this end, we first write down the kinetic energy as a functional on tangent bundle:
$L:(X,\dot{X})\in T_{X}\mathsf{SDiff}(D)\to L(X,\dot{X}):=\frac{1}{2}\int_{D}\lvert\dot{X}(t,x)\rvert^{2}\rho_{0}(x)dx$
It’s not hard to show that the action-minimizing trajectory augmented by the incompressibility constraint $\det(\nabla X)=1$ is precisely the equation (1). We refer the interested readers to [[LLP11]] for study of relaxed geodesic problem in the sense of Brenier’s generalized flow (cf. [Br89]).
Abusing the notation a little bit, we now write the Lagrangian $L$ in an equivalent form:
$L(u,X)=\frac{1}{2}\int_{D}\lvert u\rvert^{2}(X_{\sharp}\rho_{0})(x)dx$ (28)
We now state the following:
Theorem 3.1(Hamilton-Pontryagin Principle)
Fix $T>0$, assume triplet $(X,u,\xi)\in C^{1}([0,T];\mathsf{SDiff}(D)\times\mathfrak{X}_{\sigma}(D)\times\mathfrak{X}(D))$ is a critical point of the augmented action
$\Phi(u,X,\xi)=\int_{0}^{T}L(u_{t},X_{t})-\langle\xi_{t},\dot{X}-u(X)\rangle_{L_{\rho_{0}}^{2}}dt$ (29)
Then the following statements hold:
  • (I)
    $(u_{t},\rho_{0}\circ X_{t}^{-1})$ solves the incompressible inhomogeneous Euler equation (1).
  • (II)
    Consider the following Hamiltonian:
    $H(\xi,X,u)=L(u,X)+\langle\xi,u(X)\rangle_{L_{\rho_{0}}^{2}}.$
    Then $(X,u,\xi)$ solves the associated Hamiltonian equation
    $\dot{X}=\frac{\partial H}{\partial\xi},\quad\dot{\xi}=-\frac{\partial H}{\partial X},\quad\frac{\partial H}{\partial u}=0.$ (30)
  • (III)
    Given initial data $(u_{0},\xi_{0})$ satisfying the compatibility condition, the solution of (30) is given by the following Lagrangian-Eulerian system:
    $\left\{\begin{aligned} &\dot{X}=u(t,X),A=X^{-1}\\ &\rho(t,\cdot)=\rho_{0}\circ A_{t}\\ &u=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg]\end{aligned}\right$. (31)
Proof. (I) As a reminder of our notation, we denote by $A=X^{-1}$ and $\rho:=\rho_{0}\circ A$. We compute the first variation of $\Phi$. We consider smooth perturbations $\delta u\in C^{1}([0,T];\mathfrak{X}_{\sigma}(D))$, $\delta X\in C_{c}^{1}([0,T];\mathsf{SDiff}(D))$ and $\delta\xi\in C^{1}([0,T];\mathfrak{X}_{\sigma}^{*}(D))$ of $\Phi$, respectively, and compute the corresponding variational equations as follows:
  • (i)
    The variation in the variable $u$ is given by
    $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u+\varepsilon\delta u,X,\xi)=\int_{0}^{T}\int_{\mathbb{R}^{d}}\langle u\rho,\delta u\rangle-\langle\xi,\delta u(X)\rangle_{L_{\rho_{0}}^{2}}dxdt.$ (32)
    Hence vanishing of (32) implies that $\rho(u-\xi\circ A)$ is a gradient, i.e. there exists some scalar function $q$ such that
    $u\rho=(\rho_{0}\xi)\circ A+\nabla q$ (33)
  • (ii)
    The variation in variable $X$ is given by
    $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u,X+\varepsilon\delta X,\xi)$
    $\int_{0}^{T}\int_{D}\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\frac{1}{2}\lvert u\rvert^{2}\rho_{0}\circ X_{\varepsilon}^{-1}+\langle\xi,\dot{X}+\varepsilon\dot{\delta X}\rangle_{L_{\rho_{0}}^{2}}-\langle\xi,u(X+\varepsilon\delta X)\rangle_{L_{\rho_{0}}^{2}}dxdt$ (34)
    To compute $\frac{d}{d\varepsilon}X_{\varepsilon}^{-1}$, notice:
    $0=\frac{d}{d\varepsilon}(X_{\varepsilon}\circ X_{\varepsilon}^{-1})=\delta X\circ X_{\varepsilon}^{-1}+\nabla X_{\varepsilon}\circ X_{\varepsilon}^{-1}\cdot\frac{d}{d\varepsilon}X_{\varepsilon}^{-1}$
    hence we conclude:
    $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}X_{\varepsilon}^{-1}=-[(\nabla X_{\varepsilon})^{-1}\delta X]\circ X_{\varepsilon}^{-1}\bigg\rvert_{\varepsilon=0}=-[(\nabla X)^{-1}\delta X]\circ A$ (35)
    Therefore, combining (34) and (35) we conclude
    $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u,X+\varepsilon\delta X,\xi)$
    $\int_{0}^{T}\int_{D}-\frac{1}{2}\lvert u\rvert^{2}\nabla\rho_{0}\circ A\cdot[(\nabla X)^{-1}\delta X]\circ A+\langle\xi,\dot{\delta X}\rangle_{L_{\rho_{0}}^{2}}-\langle\xi,\nabla u(X)\delta X\rangle_{L_{\rho_{0}}^{2}}dxdt$
    $\int_{0}^{T}\int_{\mathbb{R}^{d}}\langle\rho_{0}\dot{\xi}+\rho_{0}\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X,\delta X\rangle dxdt.$ (36)
    Hence, sending (36), the variational equation in $X$ reads
    $\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0$ (37)
  • (iii)
    The variation in variable $\xi$ is simply given by
    $\dot{X}=u(X)$ (38)
    Hence, the critical point equation is obtained by putting together (33)(36)(38):
    $\left\{\begin{aligned} &\dot{X}=u(X),\quad A=X^{-1}\\ &\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0\\ &u\rho=(\rho_{0}\xi)\circ A+\nabla q,\quad\nabla\cdot u=0\\ &\rho=X_{*}\rho_{0}=\rho_{0}\circ A\end{aligned}\right$. (39)
Now denote by $\eta=(\rho_{0}\xi)\circ A$, by method of characteristics, (37) implies
$\partial_{t}\eta=(\rho_{0}\dot{\xi})\circ A+\nabla(\rho_{0}\xi)\circ A\cdot\partial_{t}A=-\nabla^{*}u(\rho_{0}\xi)\circ A-\frac{1}{2}\lvert u\rvert^{2}\nabla\rho-u\cdot\nabla[(\rho_{0}\xi)\circ A].$
hence $\eta$ solves the PDE:
$\partial_{t}\eta+u\cdot\nabla\eta+\nabla^{*}u\cdot\eta+\frac{1}{2}\lvert u\rvert^{2}\nabla\rho=0.$ (40)
Now, since $u\rho=\nabla q+\eta$, plugging in the equation (40) we obtain
$\partial_{t}(u\rho-\nabla q)+u\cdot\nabla(u\rho-\nabla q)+\nabla^{*}u\cdot(u\rho-\nabla q)+\frac{1}{2}\lvert u\rvert^{2}\nabla\rho=0$
Simple rearrangement yields
$\partial_{t}(\rho u)+u\cdot\nabla(\rho u)+\nabla\bigg(\frac{1}{2}\lvert u\rvert^{2}\rho-u\cdot\nabla q-\partial_{t}q\bigg)=0.$
Hence, we let
$p=\frac{1}{2}\lvert u\rvert^{2}\rho-u\cdot\nabla q-\partial_{t}q$
and conclude that the coupled system (39) is equivalent to the system (1) with pressure $p$.
(II) We compute the partial derivatives $\partial_{\xi}H$, $\partial_{X}H$. Recall the definition of variational derivative:
$\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi+\varepsilon\delta\xi,X,u)=\int_{D}[\frac{\partial H}{\partial\xi}(\xi,X,u)\cdot\delta\xi]\rho_{0}(a)da$
and
$\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi,X+\varepsilon\delta X,u)=\mathsf{G}_{X}\big(\frac{\partial H}{\partial X},\delta\xi\big)=\int_{D}[\frac{\partial H}{\partial X}(\xi,X,u)\cdot\delta X]\rho_{0}(a)da$
which implies that
$\frac{\partial H}{\partial\xi}=u(X),\quad\frac{\partial H}{\partial X}=\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u(X)\rvert^{2}(\nabla^{*}X)^{-1}\nabla\log\rho_{0}$
and
$\frac{\partial H}{\partial u}=u\rho-(\rho_{0}\xi)\circ A-\nabla q.$
Putting above together, straightforward computation yields that (30) is equivalent to (39).
(III) We consider the following general version of (37):
$\dot{w}+\nabla^{*}u(X)w=f(X),\quad w(0,\cdot)=w_{0}$
Assuming sufficient regularity on $X$, applying Duhamel’s formula we have the following Lagrangian formulae for solution $w(t,\cdot)$:
$w(t,x)=(\nabla^{*}X_{t})^{-1}\bigg(w_{0}(x)+\int_{0}^{t}(\nabla^{*}X_{\tau})f(t,X_{\tau}(x))d\tau\bigg)$
Hence, we obtain the following representation for the solution $\eta$ of (40):
$\eta(t,x)=\nabla^{*}A_{t}\bigg(\rho_{0}\xi_{0}-\frac{1}{2}\int_{0}^{t}(\nabla^{*}X_{\tau})\lvert u(\tau,X)\rvert^{2}\nabla\rho(\tau,X_{\tau}(x))d\tau\bigg)\circ A_{t}(x)$
$\nabla^{*}A_{t}\bigg(\rho_{0}\xi_{0}-\frac{1}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}\nabla\rho_{0}(x)d\tau\bigg)\circ A_{t}(x).$ (41)
And we are ready to conclude:
$u=\mathbf{P}_{\rho}(\rho^{-1}\eta)=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg]$ (42)
Combining (38) (42) and $\rho=\rho_{0}\circ A$, we recover the system (31). ∎
3.2 Vorticity Formulation
We now move on to derive a stream-vorticity formulation of (1). We will state a more general functional analytic framework and treat IIE as a specific example.
Let $\mathcal{V}$ be a Hilbert space with inner product $\langle\cdot,\cdot\rangle_{\mathcal{V}}$. Let $\mathcal{I}:\mathcal{V}\to\mathcal{V}^{*}$ be the Riesz isomorphism between dual pairing $\langle,\rangle_{*}$ of $\mathcal{V},\mathcal{V}^{*}$:
$\langle v,v^{\prime}\rangle_{*}=\langle v,\mathcal{I}^{-1}v^{\prime}\rangle_{\mathcal{V}}$
Let $\mathcal{D}:\mathcal{F}\to\mathcal{V}$ be a linear operator from Hilbert space $\mathcal{F}$ (with dual pairing $\langle,\rangle$) to the Hilbert space $\mathcal{V}$ with closed range, then we have the following pull-back inner product on $\mathcal{F}$:
$\langle\varphi,\psi\rangle_{\mathcal{F}}:=\langle\mathcal{D}\varphi,\mathcal{D}\psi\rangle_{\mathcal{V}}=\langle\mathcal{D}\varphi,\mathcal{I}\mathcal{D}\psi\rangle_{*}=\langle\varphi,\mathcal{D}^{*}\mathcal{I}\mathcal{D}\psi\rangle$ (43)
Denote by $\mathcal{A}:\mathcal{F}\to\mathcal{F}^{*}$ the Riesz isomorphism on $(\mathcal{F},\langle,\rangle_{\mathcal{F}})$ we have for $\psi^{\prime}\in\mathcal{F}^{*}$ with $\mathcal{A}\psi=\psi^{\prime}$:
$\langle\varphi,\mathcal{A}\psi\rangle=\langle\varphi,\mathcal{D}^{*}\mathcal{I}\mathcal{D}\psi\rangle$ (44)
Hence we conclude: $\mathcal{A}=\mathcal{D}^{*}\mathcal{I}\mathcal{D}$. While the projection onto the range $\mathsf{Ran}(\mathcal{D})$ is:
$\mathcal{P}=\mathcal{D}\mathcal{A}^{-1}\mathcal{D}^{*}\mathcal{I}:\mathcal{V}\to\mathsf{Ran}(\mathcal{D})$
In our hydrodynamic setting, assume $d=2$ , let $\mathcal{V}=L^{2}(D;\mathbb{R}^{d})$ be the space of $L^{2}$ vector fields, $\mathcal{F}=H^{1}(D)/\mathbb{R}$, $\mathcal{F}^{*}=H_{0}^{-1}(D)$ be the space of mean-zero $H^{-1}$ distributions. Set $\mathcal{D}=\nabla^{\perp}$, hence $\mathcal{D}^{*}=\nabla\times$. The inertia operator in this case is the following uniformly elliptic operator (due to (A)):
$\mathcal{A}=\nabla\times(\rho\nabla^{\perp})=\mathbf{L}_{\rho}=\Delta_{1/\rho}$
and the projector reads:
$\mathcal{P}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho(\cdot))$
Indeed, decomposing $\rho^{-1}w=v+\rho^{-1}\nabla p$ as in proposition 2.3 (i), we have the following
$\mathcal{P}(\rho^{-1}w)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho v)=:\mathcal{K}_{\rho}\nabla\times(\rho v)$
Since $d=2$ and $\nabla\cdot v=0$, consider the stream function $\psi$ of $v$, we find:
$\mathcal{P}(\rho^{-1}w)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho\nabla^{\perp}\psi)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\mathbf{L}_{\rho}\psi=\nabla^{\perp}\psi=v.$
Therefore $\mathcal{P}=\mathbf{P}_{\rho}$. Similar calculation shows that $\mathcal{P}=\mathbf{P}_{\rho}$ also holds in $d=3$, if we replace $\nabla^{\perp}$ with $\nabla\times$ and define $\mathbf{L}_{\rho}$ as the following weighted Laplacian for vector fields:
$\mathbf{L}_{\rho}:w\in\mathfrak{X}(D)\to\nabla\times(\rho\nabla\times w)$
We formally unify the above relations in dimension $2$ in the following commutative diagram.
$\mathcal{F}:=H^{1}(D)/\mathbb{R}$$\mathcal{V}:=L^{2}_{\rho}(D;\mathbb{R}^{2})$$\mathcal{F}^{*}:=H_{0}^{-1}(D)$$\mathcal{V}^{*}\cong L^{2}_{\rho}(D;\mathbb{R}^{2})$$\scriptstyle{\mathcal{D}=\nabla^{\perp}}$$\scriptstyle{\mathcal{I}:\ v\mapsto\rho v}$$\scriptstyle{\mathbf{P}_{\rho}=\mathcal{D}\mathcal{A}^{-1}\mathcal{D}^{*}\mathcal{I}}$$\scriptstyle{\mathcal{A}^{-1}=\Delta_{\rho}^{-1}}$$\scriptstyle{\mathcal{K}_{\rho}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}}$$\scriptstyle{\mathcal{D}^{*}=\nabla\times}$ (45)
Following the geometric picture above, we now introduce the vorticity formulation. Let $\eta=\rho u+\nabla q$, we therefore have $\mathbf{P}_{\rho}(\rho^{-1}\eta)=u$. Following our above computation, we have yet another closed-form representation of $\mathbf{P}_{\rho}$:
$\mathbf{P}_{\rho}(\rho^{-1}\eta)=\nabla^{\perp}\Delta_{\rho}^{-1}\nabla\times\eta.$
Now we define $\omega=\nabla\times\eta$, then taking the curl of equation (30) we obtain the following vorticity equation:
$\partial_{t}\omega+\mathscr{L}_{u}\omega+\nabla\frac{1}{2}\lvert u\rvert^{2}\times\nabla\rho=0.$ (46)
where $\mathscr{L}_{u}$ is the Lie derivative of a scalar/vector field in $2d/3d$ respectively. Hence, in view of the above diagram, we have the following vorticity formulation of (1), where the pressure is deleted:
$\left\{\begin{aligned} &\partial_{t}\omega+\mathscr{L}_{u}\omega+\nabla\frac{1}{2}\lvert u\rvert^{2}\times\nabla\rho=0\\ &u=\mathcal{K}_{\rho}\omega:=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\omega\\ &\partial_{t}\rho+u\cdot\nabla\rho=0\end{aligned}\right$. (47)
We now show the conservation of vorticity (Kelvin circulation theorem) in the framework of IIE. In hydrodynamics, conservation laws coming from particle relabeling symmetry (Casimir Invariants in geometry) are known to be the consequence of duality between Lie transport (Lie elements of Lagrangian loops) and dual Lie transport (co-adjoint motion of momentum) ([HMR98][MP25]). Here, for the inhomogeneous fluid, we do not have a full conservation of circulation, due to the commutator forcing term in vortex dynamics. However, the following result holds in $d=2$:
Theorem 3.2
Assume $(u,\rho)$ is a smooth solution of IIE, then for any regular value $\alpha>0$ of $\rho_{0}$, where $E_{\alpha}:=\{\rho_{0}=\alpha\}$ is a smooth loop. Let $D_{\alpha}:=\{\rho_{0}<\alpha\}$ is a smooth bounded domain enclosed by $E_{\alpha}$, we have
$\oint_{X_{t}(E_{\alpha})}u_{t}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell.$ (48)
which implies
$\int_{X_{t}(D_{\alpha})}\omega_{t}(x)dx=\int_{D_{\alpha}}\omega_{0}(x)dx.$ (49)
Proof. Thanks to the Lagrangian formulae (41), we have for any smooth loop $\gamma$:
$\oint_{X_{t}(\gamma)}\eta_{t}\cdot d\ell=\oint_{\gamma}\nabla^{*}X_{t}\eta_{t}\circ X_{t}\cdot d\ell$
$\oint_{\gamma}\big[u_{0}\rho_{0}+\nabla q_{0}-\nabla\rho_{0}(x)\frac{1}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\big]\cdot d\ell$
$\oint_{\gamma}u_{0}\rho_{0}\cdot d\ell-\frac{1}{2}\int_{0}^{t}\oint_{\gamma}\lvert\dot{X}_{\tau}\rvert^{2}\nabla\rho_{0}\cdot d\ell d\tau.$
Now choose $\gamma=E_{\alpha}$, since $\{\rho_{t}=\alpha\}=X_{t}(E_{\alpha})$, we conclude:
$\oint_{E_{\alpha}}u_{t}\cdot d\ell=\alpha^{-1}\oint_{E_{\alpha}}\eta_{t}\cdot d\ell=\alpha^{-1}\oint_{E_{\alpha}}\rho_{0}u_{0}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell.$
where we use the orthogonality relation: For all smooth $f$ we have
$\oint_{E_{\alpha}}f\nabla\rho_{0}\cdot d\ell=\int f\nabla\rho_{0}\cdot\frac{\nabla^{\perp}\rho_{0}}{\lvert\nabla\rho_{0}\rvert}\mathscr{H}_{E_{\alpha}}^{1}(dx)=0.$
Hence, the proof of (48) is complete. To show (49), simply apply the Stokes theorem. ∎
3.3 Local Existence Via Vorticity Formulation
We recall that in the general dimension $d$, IIE is shown to be locally well-posed in $H^{s+1}\times H^{s+1}$ for $s>d/2$, and we refer the readers to [Da06][[Da10]] [[DF11]] for various well-posedness results in $L^{p}$-based Sobolev spaces and endpoint Besov spaces. Here we show the local existence in Sobolev space of the vorticity formulation (47) of the inhomogeneous Euler equation.
We restrict ourselves to the torus case $D=\mathbb{T}^{2}$ for convenience. We’ll use the notation $H_{0}^{k}$ for space of mean-zero $H^{k}(D)$ functions, and we define the weighted space:
$\dot{H}_{\rho}^{1}:=\{f:\lVert\nabla f\rVert_{L_{\rho}^{2}}<\infty\}$
We start with the following elliptic estimate.
Lemma 3.3
Assume $\omega\in H_{0}^{2}$, $\rho\in H^{3}$ satisfying the assumption (A), then any mean-zero strong solution $\psi$ of $\mathbf{L}_{\rho}\psi=\omega$ also satisfies $\psi\in H_{0}^{4}$.
Proof: We’ll use a standard bootstrap argument. The proof is divided into $3$ steps.
Step 1.($H^{1}$ Estimate) Multiplying both sides of the equation by $\psi$ and integrating by part, we have:
$\lVert\nabla\psi\rVert_{L_{\rho}^{2}}^{2}=\int_{D}\lvert\nabla\psi(x)\rvert^{2}\rho(x)dx=-\int_{D}\psi(x)\cdot(\rho^{-1}\omega)(x)\rho(x)dx\leq\lVert\psi\rVert_{L_{\rho}^{2}}\lVert\rho^{-1}\omega\rVert_{L_{\rho}^{2}}.$
Now we control the $L_{\rho}^{2}$ norm of $\psi$ via $L_{\rho}^{2}$ norm of $\nabla\psi$ via Poincaré inequality. Thanks to assumption $A$ and Poincaré inequality, we have:
$\lVert\psi\rVert_{L_{\rho}^{2}}\leq C_{1}\lVert\psi\rVert_{L^{2}}\leq C_{1}C_{P}\lVert\nabla\psi\rVert_{2}\leq C_{1}^{2}C_{P}\lVert\nabla\psi\rVert_{L_{\rho}^{2}}.$
Where $C_{P}$ is the bounding constant of Poincaré inequality. We therefore have:
$C_{1}^{-2}\lVert\nabla\psi\rVert_{L^{2}}^{2}\leq\lVert\nabla\psi\rVert_{L_{\rho}^{2}}^{2}\leq C_{1}^{2}C_{P}\int_{D}\frac{\omega^{2}}{\rho}dx\leq C_{1}^{3}C_{P}\lVert\omega\rVert_{L^{2}}^{2}.$
Hence, we conclude $\lVert\nabla\psi\rVert_{2}\lesssim\lVert\omega\rVert_{2}$.
Step 2.($H^{2}$ Estimate) Now we expand the equation and normalize the leading coefficient to obtain:
$\Delta\psi+\nabla\psi\cdot\nabla\log\rho=\rho^{-1}\omega$ (50)
Simple rearrangement of (50) yields:
$\lVert\Delta\psi\rVert_{2}\leq\lVert\rho^{-1}\omega\rVert_{2}+\lVert\rho^{-1}\nabla\rho\cdot\nabla\psi\rVert_{2}\leq C_{1}\lVert\omega\rVert_{2}+C_{1}^{6}C_{P}\lVert\nabla\rho\rVert_{\infty}\lVert\omega\rVert_{2}.$
As a consequence of Sobolev embedding $H^{3}\hookrightarrow W^{1,\infty}$ in $2$-d, we have:
$\lVert\psi\rVert_{H^{2}}\lesssim(1+\lVert\rho\rVert_{H^{3}})\lVert\omega\rVert_{2}$
Step 3. (Bootstrap to Higher Order) Now we bootstrap the regularity in (50). A higher-order estimate yields:
$\lVert\Delta\psi\rVert_{H^{1}}\leq\lVert\rho^{-1}\omega\rVert_{H^{1}}+\lVert\rho^{-1}\nabla\rho\cdot\nabla\psi\rVert_{H^{1}}\lesssim\lVert\rho^{-1}\rVert_{H^{2}}\lVert\omega\rVert_{H^{1}}+\lVert\rho^{-1}\rVert_{H^{2}}\lVert\nabla\rho\rVert_{H^{2}}\lVert\psi\rVert_{H^{2}}$
where we use the algebra property of $H^{2}$ and the following para-product inequality (cf. [Wi22]):
$\lVert fg\rVert_{H^{1}}\leq\lVert f\rVert_{H^{1}}\lVert g\rVert_{H^{r}},\quad r>1$ (51)
Now, since $\rho\in H^{3}$ and the function $f(z)=z^{-1}$ is smooth on the range of $\rho$, we have $\rho^{-1}=f(\rho)\in H^{3}$. Hence, we conclude:
$\lVert\psi\rVert_{H^{3}}\lesssim\lVert\omega\rVert_{H^{1}}\lVert\rho^{-1}\rVert_{H^{2}}(1+\lVert\rho\rVert_{H^{3}}+\lVert\rho\rVert_{H^{3}}^{2})\lesssim\lVert\omega\rVert_{H^{1}}\sum_{j=1}^{3}\lVert\rho\rVert_{H^{3}}^{j}$
Finally, plugging the above estimate back to (50), we conclude
$\lVert\psi\rVert_{H^{4}}\lesssim\lVert\omega\rVert_{H^{2}}P(\lVert\rho\rVert_{H^{3}})$ (52)
where $P$ is some polynomial of degree $5$ with positive coefficients, hence it is monotonically increasing on $(0,\infty)$. ∎
We now prove the local existence of IIE using our vorticity formulation.
Theorem 3.4
Let $(\omega_{0},\rho_{0})\in H^{2}\times H^{3}$ and let $\rho$ satisfies Assumption (A), then (47) admits a unique solution local in time.
Proof. We consider the following iterative scheme:
$\left\{\begin{aligned} &\partial_{t}\omega^{n+1}+{u^{n}}\cdot\nabla\omega^{n+1}+\nabla\frac{1}{2}\lvert u^{n}\rvert^{2}\times\nabla\rho^{n}=0\\ &\partial_{t}\rho^{n+1}+u^{n}\cdot\nabla\rho^{n+1}=0\\ &u^{n+1}=\mathcal{K}_{\rho}^{n+1}\omega^{n+1}:=\nabla^{\perp}\mathbf{L}_{\rho^{n+1}}^{-1}\omega^{n+1}\end{aligned}\right$. (53)
Step 1. We define
$\mathcal{U}_{R}=\{(\omega,u,\rho)\in C_{t}H^{2}\times C_{t}H_{\sigma}^{3}\times C_{t}H^{3}:\lVert\omega\rVert_{C_{t}H_{x}^{2}}\leq R_{1},\lVert u\rVert_{C_{t}H_{x}^{3}}\leq R_{2},\lVert\rho\rVert_{C_{t}H_{x}^{3}}\leq R_{3}\}$
We aim to show that for sufficiently small time $T$ and positive numbers $R_{1},R_{2},R_{3}$ that we’ll choose later, the sequence $\{(\omega^{n},u^{n},\rho^{n})\}$ generated by the iterative scheme (1) stays in $U_{R}$.
First, notice that:
$\rho^{n+1}(t,\cdot)=\rho_{0}\circ A_{t}^{n},\quad\text{where }\dot{X}_{t}^{n}=u_{t}^{n}(X_{t}^{n}),\quad A_{t}^{n}=(X_{t}^{n})^{-1}$
and $1/\rho^{n+1}$ is also transported. Hence, we have
$\lVert\rho_{t}^{n+1}\rVert_{H^{3}},\lVert 1/\rho_{t}^{n+1}\rVert_{H^{3}}\leq C(\rho_{0})\exp\bigg(C\int_{0}^{t}\lVert u^{n}\rVert_{H^{3}}d\tau\bigg)\leq C(\rho_{0})\exp(CtR_{2}).$
Here, $C(\rho_{0})$ can be chosen as $C(\rho_{0})=\lVert\rho_{0}\rVert_{H^{3}}+\lVert\rho_{0}^{-1}\rVert_{H^{3}}$.
Now we apply $J^{2}:=1-\Delta$ to the vorticity equation, pairing with $J^{2}\omega^{n}$ and integrating by part we have:
$\frac{d}{dt}\frac{1}{2}\lVert\omega^{n+1}\rVert_{H^{2}}^{2}=\langle\partial_{t}J^{2}\omega^{n+1},J^{2}\omega^{n+1}\rangle_{2}$
$\langle J^{2}(u^{n}\cdot\nabla\omega^{n+1})-u^{n}\cdot\nabla J^{2}\omega^{n+1},J^{2}\omega^{n+1}\rangle_{2}-\langle u^{n}\cdot\nabla J^{2}\omega^{n+1},J^{2}\omega^{n+1}\rangle_{2}+\langle J^{2}(\{\frac{1}{2}\lvert u^{n}\rvert^{2},\rho^{n}\}),J^{2}\omega^{n+1}\rangle_{2}$
Apply Kato-Ponce commutator estimate (cf. [KP88]) on the first term, integrating the second term by part and deleting it using divergence-free property of $u^{n}$, and using the algebra property of $H^{2}$ we conclude
$\frac{d}{dt}\frac{1}{2}\lVert\omega^{n+1}\rVert_{H^{2}}^{2}\lesssim\lVert\nabla u^{n}\rVert_{H^{2}}\lVert\omega^{n+1}\rVert_{H^{2}}^{2}+\lVert u^{n}\rVert_{H^{3}}\lVert\rho^{n}\rVert_{H^{2}}\lVert\omega^{n+1}\rVert_{H^{2}}$
which, together with our assumption $(\omega^{n},u^{n},\rho^{n})\in\mathcal{U}_{R}$, gives:
$\frac{d}{dt}\lVert\omega^{n+1}\rVert_{H^{2}}\lesssim R_{2}\lVert\omega^{n+1}\rVert_{H^{2}}+R_{2}R_{3}$
Gronwall inequality therefore implies:
$\lVert\omega^{n+1}(t)\rVert_{H^{2}}\leq\exp(CR_{2}t)(\lVert\omega_{0}\rVert_{H^{2}}+CR_{2}R_{3}t)$
Here we temporarily choose $T$ to be sufficiently small such that
${CR_{2}T}\leq 1/2,\quad\frac{1}{2}\sqrt{e}R_{3}\leq\lVert\omega_{0}\rVert_{H^{2}}$
where $C>1$ is some generic large number. Later we’ll fix $R_{2}$. Thanks to this choice, for all $t\in[0,T]$ we have:
$\lVert\rho^{n+1}\rVert_{H^{3}},\lVert 1/\rho^{n+1}\rVert_{H^{3}}\leq\sqrt{e}C(\rho_{0}),\quad\lVert\omega_{t}^{n+1}\rVert_{H^{2}}\leq 2\sqrt{e}\lVert\omega_{0}\rVert_{H^{2}}$
Hence we can choose $R_{1}=2\sqrt{e}\lVert\omega_{0}\rVert_{H^{2}}$.
Now we estimate $u^{n+1}$. Given $\lVert\omega^{n+1}\rVert\leq 2\sqrt{e}\lVert\omega_{0}\rVert_{H^{2}}$ and $\lVert\rho^{n+1}\rVert_{H^{3}}\leq\sqrt{e}C(\rho_{0})$, we have:
$\lVert u^{n+1}\rVert_{H^{3}}\lesssim\lVert\mathbf{L}_{\rho^{n+1}}^{-1}\omega^{n+1}\rVert_{H^{4}}=\lVert\psi^{n+1}\rVert_{H^{4}}\leq 2\sqrt{e}CP(\sqrt{e}C(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}}$
where $P$ is the same polynomial as in (52). Hence, we may choose:
$R_{2}=2\sqrt{e}CP(\sqrt{e}C(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}},\quad R_{3}=\sqrt{e}C(\rho_{0})$
And the above analysis demonstrates that our iteration scheme stays in $\mathcal{U}_{R}$.
Step 2. Now, we show that the above iterative sequence $\{(\omega^{n},u^{n},\rho^{n})\}\subset\mathcal{U}_{R}$ is Cauchy in $C_{t}H^{1}\times C_{t}H_{\sigma}^{2}\times C_{t}H^{2}$. We start by noticing for $u^{n}=\nabla^{\perp}\psi^{n}$:
$\frac{\mathbf{L}_{\rho^{n+1}}\psi^{n+1}}{\rho^{n+1}}-\frac{\mathbf{L}_{\rho^{n}}\psi^{n}}{\rho^{n}}=\frac{\omega^{n+1}}{\rho^{n+1}}-\frac{\omega^{n}}{\rho^{n}}$
Which implies:
$\Delta(\psi^{n+1}-\psi^{n})+\nabla\log\rho^{n+1}\cdot\nabla(\psi^{n+1}-\psi^{n})+\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}=\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}+(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}$
Denote by $C_{R}$ constants depending on $R_{1},R_{2},R_{3}$. We now again make use of the (51) to control the RHS:
$\bigg\lVert\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\omega^{n+1}-\omega^{n}\rVert_{H^{1}},\quad\bigg\lVert(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}}$
and the commutator term
$\lVert\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}}$
Hence, following the same procedure in lemma 3.3, let $\delta\omega^{n}:=\omega^{n+1}-\omega^{n}$, $\delta\rho^{n}:=\rho^{n+1}-\rho^{n}$ and $\delta u^{n}=u^{n+1}-u^{n}$, we conclude
$\lVert\delta u^{n}\rVert_{H^{2}}\leq\lVert\psi^{n+1}-\psi^{n}\rVert_{H^{3}}\leq C_{R}(\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}})$ (54)
Straightforward computation yields the following equation for $\delta\omega^{n}$ and $\delta\rho^{n}$:
$\partial_{t}\delta\omega^{n}+u^{n}\cdot\nabla\delta\omega^{n}+\delta u^{n}\cdot\nabla\omega^{n}=\frac{1}{2}(\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\}):=\sigma_{n}$
$\partial_{t}\delta\rho^{n}+u^{n}\cdot\nabla\delta\rho^{n}+\delta u^{n}\cdot\nabla\rho^{n}=0$
Again, applying $J$, $J^{2}$ to above two equations, pairing with $J\delta\omega^{n}$, $J^{2}\delta\rho^{n}$ respectively, thanks to div-free property of $u^{n}$ we have
$\frac{d}{dt}\frac{1}{2}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+\langle J(u^{n}\cdot\nabla\delta\omega^{n})-u^{n}\cdot\nabla J\delta\omega^{n},J\delta\omega^{n}\rangle_{2}+\langle J\nabla\cdot(\delta u^{n}\omega^{n}),J\delta\omega^{n}\rangle_{2}=\langle J\sigma^{n},J\delta\omega^{n}\rangle$
$\frac{d}{dt}\frac{1}{2}\lVert\delta\rho^{n}\rVert_{H^{2}}^{2}+\langle J^{2}(u^{n}\cdot\nabla\delta\rho^{n})-u^{n}\cdot\nabla J^{2}\delta\rho^{n},J^{2}\delta\rho^{n}\rangle_{2}+\langle J^{2}(\delta u^{n}\cdot\nabla\rho^{n}),J^{2}\delta\rho^{n}\rangle_{2}=0$
Now, thanks to Kato-Ponce and the algebra property of $H^{2}$, $H^{3}$, we conclude
$\frac{d}{dt}\frac{1}{2}\lVert\delta\rho^{n}\rVert_{H^{2}}^{2}\leq C_{R}\lVert\delta\rho^{n}\rVert_{H^{2}}^{2}+C_{R}\lVert\delta u^{n}\rVert_{H^{2}}\lVert\delta\rho^{n}\rVert_{H^{2}}$ (55)
$\frac{d}{dt}\frac{1}{2}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}\leq C_{R}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+C_{R}\lVert\delta u^{n}\rVert_{H^{2}}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\sigma^{n}\rVert_{H^{1}}\lVert\delta\omega^{n}\rVert_{H^{1}}$
Notice that
$\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\}=\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n}\rvert^{2},\rho^{n-1}\}+\{\lvert u^{n}\rvert^{2},\rho^{n-1}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\}$
$\{\lvert u^{n}\rvert^{2},\delta\rho^{n-1}\}+\nabla^{*}u^{n}(u^{n}-u^{n-1})\nabla^{\perp}\rho^{n-1}+\nabla^{*}(u^{n}-u^{n-1})u^{n-1}\nabla^{\perp}\rho^{n-1}$
$\{\lvert u^{n}\rvert^{2},\delta\rho^{n-1}\}+\nabla^{*}u^{n}\delta u^{n-1}\nabla^{\perp}\rho^{n-1}+\nabla^{*}\delta u^{n-1}u^{n-1}\nabla^{\perp}\rho^{n-1}$
Hence, we conclude that
$\frac{d}{dt}\frac{1}{2}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}\leq C_{R}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+C_{R}\lVert\delta u^{n}\rVert_{H^{2}}\lVert\delta\omega^{n}\rVert_{H^{1}}+C_{R}(\lVert\delta\rho^{n-1}\rVert_{H^{2}}+\lVert\delta u^{n-1}\rVert_{H^{2}})\lVert\delta\omega^{n}\rVert_{H^{1}}$ (56)
Step 3. Now we define: $E^{n}(t):=\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+\lVert\delta\rho^{n}\rVert_{H^{1}}^{2}$. By (54), (55), (56), we obtain:
$\frac{d}{dt}E^{n}(t)\leq C_{R}E^{n}(t)+C_{R}E^{n-1}(t),\quad E(0)=0$
Now, integrating above inequality from time 0 to $t$, and denote by $U_{n}:=\sup_{t\in[0,T]}E^{n}(t)$. We have:
$E^{n}(t)\leq C_{R}\int_{0}^{t}E^{n}(\tau)d\tau+C_{R}\int_{0}^{t}E^{n-1}(\tau)d\tau\leq C_{R}TU_{n}+C_{R}TU_{n-1}.$
Therefore, taking supremum over $t$ on the LHS we conclude:
$U_{n}\leq C_{R}TU_{n}+C_{R}TU_{n-1}.$
Choosing $T$ smaller if necessary, such that $C_{R}T\leq 1/3$, we have
$\frac{2}{3}U_{n}\leq(1-C_{R}T)U_{n}\leq C_{R}TU_{n-1}\leq\frac{1}{3}U_{n-1}.$
Therefore, we obtain $U_{n}\leq U_{n-1}/2\leq 2^{-(n-1)}U_{1}$. In particular, we demonstrate summability of sequence
$U_{n}=\sup_{t\in[0,T]}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+\lVert\delta\rho^{n}\rVert_{H^{1}}^{2}.$
As a matter of fact, we conclude that $(\omega^{n},\rho^{n})$ is Cauchy in $C([0,T];H^{1}\times H^{2})$. Now, $u^{n}$ is also Cauchy in $C([0,T];H_{\sigma}^{2})$, thanks to (54). We therefore have a unique limit $(\omega,u,\rho)$.
Finally, closedness of $\mathcal{U}_{R}$ in $C([0,T];H^{1}\times H_{\sigma}^{2}\times H^{2})$ guarantees that $(\omega,u,\rho)\in\mathcal{U}_{R}$, which is the unique solution of (47). ∎
4 Lagrangian Analyticity of Solution
Lagrangian trajectories of many incompressible hydrodynamic models are known to be locally analytic in time [[FZ14]][CVW15], as long as the initial data lies in the local well-posedness class. On the other hand, such a nice property fails to hold when we switch to the compressible models (cf. [He19], since the pressure is no longer an elliptic constraint, but forces the system to be hyperbolic conservation law with finite speed of propagation. Therefore, there exists points which stays rest and then starts to move, prevents analyticity of Lagrangian displacement.
In this section, we approach the Lagrangian analyticity of inhomogeneous incompressible Euler equation using our Lagrangian formulation. Our approach is similar to [[FZ14]], where the vorticity formulation is used and Cauchy’s invariant plays a key role towards the iterative scheme for Taylor coefficients. In the present work, we stick to the momentum level, which allows us to deal with general dimensions in a unified way. The main theorem of this section is as follows.
Theorem 4.1
Let $s>d/2$ and assume $u_{0}\in H_{\sigma}^{s+1}$, $\rho_{0}\in H^{s+1}$ satisfies assumption (A), then there exists a time threshold $T_{*}$ such that the Lagrangian flow $X$ of solution $u$ of IIE is analytic for $t\in(0,T_{*})$
It turns out that inhomogeneity doesn’t prevent analyticity, by comparing above theorem 4.1 with the compressible Euler counterexample in [He19]. The crucial structure of incompressible fluids that leads to the analyticity is the hidden ellipticity in these models (The Biot-Savart law). Such a regularization mechanism, from the Lagrangian viewpoint, is manifested by the following pointwise integration by part lemma (cf. [Co01a][CI08]).
Lemma 4.2
Let $\ell,u\in H^{s+1}$, $s>d/2$. Then $\mathbf{P}(\nabla^{*}\ell u)\in H^{s+1}$.
Proof. It suffices to show $\partial_{i}\mathbf{P}(\nabla^{*}\ell u)\in H^{s}$ for all $i$. For each fixed $1\leq i\leq d$, by commutativity between Fourier multipliers we have
$\partial_{i}\mathbf{P}(\partial_{k}\ell_{j}u_{j})=\mathbf{P}(\partial_{ki}\ell u_{j}+\partial_{k}\ell_{j}\partial_{i}u_{j})=\mathbf{P}(\partial_{k}(\partial_{i}\ell_{j}u_{j})-\partial_{i}\ell_{j}\partial_{k}u_{j}+\partial_{k}\ell_{j}\partial_{i}u_{j})$
$\mathbf{P}(-\partial_{i}\ell_{j}\partial_{k}u_{j}+\partial_{k}\ell_{j}\partial_{i}u_{j})\in H^{s}$
where the last line follows from the fact that $H^{s}$ is an algebra and $\mathbf{P}$ is a 0-order multiplier. ∎
Proof of Theorem 4.1 We’ll use our Lagrangian representation to obtain a recursive formula for Taylor coefficients of Lagrangian displacement, and prove convergence of the Taylor series. The proof is divided into steps.
Step 1.(Set up the Lagrangian Framework)
Recall that in proof of theorem 3.1 (III) we derived:
$\eta_{t}=\rho_{t}u_{t}+\nabla q_{t}=\nabla^{*}A_{t}\bigg(\eta_{0}-\int_{0}^{t}\frac{1}{2}\lvert\dot{X}_{\tau}(x)\rvert^{2}\nabla\rho_{0}(x)d\tau\bigg)\circ A_{t}$
Acting $X_{t}$ to the above equality in the sense of pull-back of covector we have
$\nabla^{*}X_{t}(\rho_{t}u_{t}+\nabla q_{t})\circ X_{t}=\eta_{0}-\nabla\rho_{0}\int_{0}^{t}\frac{1}{2}\lvert\dot{X}_{\tau}\rvert^{2}d\tau$
Denote by $\ell_{t}(a)=X_{t}(a)-a$ the Lagrangian displacement, simplify the LHS, we conclude:
$\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\nabla(q_{t}\circ X_{t})+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert(\tau)d\tau=\eta_{0}$ (57)
Which suggests that
$\frac{d^{n}}{dt^{n}}\mathbf{P}\bigg[\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert^{2}(\tau)d\tau\bigg]=0,\quad\text{for all }n\geq 1$ (58)
Meanwhile, the following determinant identity relation automatically holds:
$\det(\mathbf{Id}+\nabla\ell(t))=1\quad\text{for all }t\geq 0$ (59)
Now, thanks to (58), assuming for sufficiently small $t>0$, we have the following series expansion of $\ell$:
$\ell_{j}(t,a)=\sum_{k=1}^{\infty}t^{k}\ell_{j}^{(k)}(a)$
Notice here $\ell^{(0)}=X(0,a)-a=0$, $\ell^{(1)}=u_{0}(a)$. Now straightforward computation yields:
$(\nabla^{*}\ell(t,a))_{ij}=\sum_{k=1}^{\infty}t^{k}\partial_{i}\ell_{j}^{(k)}(a),\quad\dot{\ell}_{j}(t,a)=\sum_{k=0}^{\infty}(k+1)t^{k}\ell_{j}^{(k+1)}$
Hence we moreover have:
$(\nabla^{*}\ell\dot{\ell})_{i}=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}(m-k+1)\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m+1-k)}$
$\int_{0}^{t}\lvert\dot{\ell}(\tau)\rvert^{2}d\tau=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}\frac{k(m+1-k)}{m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}$
Combining the above together, we conclude that (58) reads for all $n\geq 1$:
$\frac{d^{n}}{dt^{n}}\bigg\lvert_{t=0}\mathbf{P}\bigg[\rho_{0}\ell_{i}^{(1)}+\sum_{m=1}^{\infty}t^{m}\bigg((m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg)\bigg]=0$
Which implies that coefficient of $t^{m}$ vanishes at all order: For all $m\geq 1$
$\mathbf{P}\bigg[(m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg]=0$ (60)
Moreover determinant identity (59) naturally implies
$\frac{d^{m}}{dt^{m}}\bigg\rvert_{t=0}\det\big(\delta_{ij}+\sum_{m=1}^{\infty}\partial_{j}\ell_{i}^{(m)}t^{m}\big)=0$
Using standard matrix identity, we have
$0=\log\det(\mathbf{Id}+\nabla\ell(t))=\mathsf{Tr}\log(\mathbf{Id}+\nabla\ell(t))=\mathsf{Tr}\bigg(\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}[\nabla\ell(t)]^{k}\bigg)$
$\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\mathsf{Tr}\bigg[\bigg(\sum_{r=1}^{\infty}\nabla\ell^{(r)}t^{r}\bigg)^{k}\bigg]$
Hence the vanishing of coefficient for $t^{m}$ implies:
$0=\sum_{k=1}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})})$ (61)
Step 2. We now derive the iterative scheme for coefficients $\{\ell^{(k)}\}_{k\geq 1}$ based on (60)(61). To this end, we rewrite (60) as
$(m+1)\ell_{i}^{(m+1)}\rho_{0}+U_{i}^{(m)}=\nabla p_{i}^{(m)}$ (62)
where $U^{(m)}$ is a vector field quadratic in lower-order terms:
$U^{(m)}=\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)$
Meanwhile, (61) implies:
$\mathsf{Tr}(\nabla\ell^{(m)})=\nabla\cdot\ell^{(m)}=\sum_{k=2}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m,r_{i}\geq 1}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})}):=W^{(m)}$ (63)
Here we remark that the expansion stops at $m=d$, eventually leading to a polynomial of $\nabla\ell^{(r)}$ with degree $d$.
Therefore, the following recursive relation holds: Fix $m$ and assume we already know $\{\ell^{(k)}\}_{k=1}^{m}$, then we can explicitly compute $U^{(m)}$, $W^{(m+1)}$ and $\nabla\cdot\ell^{(m+1)}=W^{(m+1)}$. Taking into account (62) we have:
$\nabla\cdot\ell^{(m+1)}=\frac{1}{m+1}\nabla\cdot(\rho_{0}^{-1}(\nabla p^{(m)}-U^{(m)}))=W^{(m+1)}$
which guarantees:
$\nabla p^{(m)}=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}+\nabla\cdot(\rho_{0}^{-1}U^{(m)})]$ (64)
Therefore one can produce $\ell^{(m+1)}$ from $\{\ell^{(k)}\}_{k=1}^{m}$ as follows:
$\ell^{(m+1)}=\frac{1}{(m+1)\rho_{0}}\big[\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}U^{(m)})]-U^{(m)}\big]$
Which is an explicit recursion formula for Taylor coefficients.
Step 3.(Elliptic Estimates) Introduce the following generating function:
$\bar{\zeta}(t):=\sum_{k=1}^{\infty}t^{k}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}$ (65)
We aim to show that the formal series (65) converges absolutely.
Here, for later convenience, we decompose $U^{(m)}=\mathbf{P}U^{(m)}+\nabla\varphi^{(m)}$. Notice then
$\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}U^{(m)})=\nabla\varphi^{(m)}+\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})$
First, we need to show that $\ell^{(m+1)}\in H^{s+1}$, given $\{\ell^{(k)}\}_{k=1}^{m}\subset H^{s+1}$. Note that with our assumption (A), $\Delta_{\rho_{0}}$ is uniformly elliptic. The above equality (64) reads
$\nabla(p^{(m)}-\varphi^{(m)})=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})]$ (66)
and we have by (62):
$\ell^{(m+1)}=\frac{1}{(m+1)\rho_{0}}\big[-\mathbf{P}U^{(m)}+\nabla(p^{(m)}-\varphi^{(m)})\big]$ (67)
We first control $\mathbf{P}U^{(m)}$. We notice that:
$\mathbf{P}U^{(m)}=\sum_{k=1}^{m}(m-1+k)\big(\mathbf{P}[\nabla^{*}\ell^{(k)}(\ell^{(m+1-k)}\rho_{0})]+\frac{k}{2m}\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})])$
Therefore, $\mathbf{P}U^{(m)}\in H^{s+1}$ will follow if we succeed in showing for any $1\leq k\leq m$:
$\mathbf{P}[\nabla^{*}\ell^{(k)}(\ell^{(m+1-k)}\rho_{0})]\in H^{s+1},\quad\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]\in H^{s+1}$
Following from combining Lemma 4.2 and a simple manipulation on the second term:
$\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]=-\mathbf{P}[\nabla(\ell^{(m+1-k)}\cdot\ell^{(k)})\rho_{0}]=-\mathbf{P}[\nabla^{*}\ell^{(m+1-k)}\cdot\ell^{(k)}\rho_{0}]-\mathbf{P}[\nabla^{*}\ell^{(k)}\cdot\ell^{(m+1-k)}\rho_{0}]$
We eventually conclude the following estimate:
$\lVert\mathbf{P}U^{(m)}\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}$ (68)
On the other hand, rewrite (64) as
$\nabla(p^{(m)}-\varphi^{(m)})=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})]$
Now, a slight variant of lemma 3.3 (taking $\rho=1/\rho_{0}$) implies:
$\lVert\nabla(p^{(m)}-\varphi^{(m)})\rVert_{H^{s+1}}\leq C(\rho_{0})(m+1)\lVert W^{(m+1)}\rVert_{H^{s}}+\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}$
Since $W^{(m+1)}$ consists of the sum of product of $\nabla\ell^{(r)}$ with $1\leq r\leq m$, algebra property of $H^{s}$ implies $W^{(m+1)}\in H^{s}$, with:
$\lVert W^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}$ (69)
while above estimate (68) implies
$\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}$
Hence we conclude from (67):
$\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}+\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}$ (70)
where RHS is a polynomial of $\lVert\nabla\ell^{(r)}\rVert_{H^{s}}$ up to order $d$. Hence, there exists positive numbers $C_{2},\cdots,C_{d}$ such that
$\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\leq\sum_{j=2}^{d}C_{j}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}$ (71)
Step 4. (Conclude Analyticity) Now based on the recurrence relation, we can recursively build a sequence $\{\beta_{m}\}$ dominating coefficients $\{\lVert\nabla\ell^{(m)}\rVert_{H^{s}}\}$ of our generating function $\zeta$ as follows:
$\beta_{1}=\lVert\nabla u_{0}\rVert_{H^{s}},\quad\beta_{m+1}=\sum_{j=2}^{d}C_{j}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\beta_{r_{i}}$
and we define the formal series
$\tilde{\zeta}(t)=\sum_{m=1}^{\infty}\beta_{m}t^{m}$
which is the formal expansion of the unique root of the following polynomial:
$P(t,\zeta)=-\zeta+t\lVert\nabla u_{0}\rVert_{H^{s}}+C_{2}\zeta^{2}+\cdots+C_{d}\zeta^{d}$ (72)
To see the uniqueness of root of $P(t,\cdot)$, notice first that for each $t\geq 0$, solving $P(t,\zeta)=0$ is equivalent to solve
$J_{t}(\alpha):=\alpha-\sum_{j=2}^{d}C_{j}\alpha^{j}=t\|\nabla u_{0}\rVert_{H^{s}}$
We observe that for any $t>0$, $J_{t}(0)=0$, while $J_{t}^{\prime}(0)=1>0$ and $J_{t}(\alpha)\to$-$\infty$ as $\alpha\to\infty$. Hence, we consider the first $\alpha_{*}$ with sufficiently small derivative: Fix $0<\delta\ll 1$, we consider
$\alpha^{*}=\min\{\alpha:J^{\prime}(\alpha)=\delta\},\quad J^{\prime}(\alpha)>\delta\text{ for all }0<\alpha<\alpha^{*}$
We now define
$\Psi_{t}(\zeta):=t\rVert\nabla u_{0}\rVert_{H^{s}}+\sum_{j=2}^{d}C_{j}\zeta^{j}$
Notice that
$\lvert\Psi_{t}^{\prime}(\zeta)\rvert\leq\sum_{j=2}^{d}jC_{j}\zeta^{d-1}=1-J^{\prime}(\zeta)\leq 1-\delta,\quad\text{for all }\zeta\in[0,\alpha_{*}]$
Hence for each $t$, $\Psi_{t}(\zeta)$ is a contraction. Moreover, we notice that $\Psi_{t}$ is increasing, hence:
$\Psi_{t}([0,\alpha])\subset[\Psi_{t}(0),\Psi_{t}(\alpha)]=\bigg[t\|\nabla u_{0}\rVert_{H^{s}},t\|\nabla u_{0}\rVert_{H^{s}}+\sum_{j=2}^{d}C_{j}\alpha^{j}\bigg]$
To apply Banach fixed point theorem, it suffices to show:
$t\|\nabla u_{0}\rVert_{H^{s}}+\sum_{j=2}^{d}C_{j}\alpha^{j}\leq\alpha$
Hence it suffices to pick $t$ such that
$t\leq T(\alpha):=\|\nabla u_{0}\rVert_{H^{s}}^{-1}J(\alpha)$
Therefore, for all $t\in[0,T(\alpha)]$ with $\alpha\in(0,\alpha^{*})$, $\Psi_{t}$ has a unique fixed point. Hence, we find a unique $\zeta_{*}\in(0,\alpha^{*})$ such that $P(t,\zeta^{*})=0$.
Now, since $P(t,\zeta)$ satisfies
$P(0,0)=0,\quad\partial_{w}P(0,0)=-1\neq 0$
The implicit function theorem implies that locally around $(0,0)$, there exists an analytic function $\zeta_{*}(t)$ such that $P(\zeta_{*}(t),t)=0$. By uniqueness of solution of $P(t,\cdot)$, we conclude the formal series solution must coincide with analytic function $\zeta_{*}(t)$, which concludes the absolute convergence of series $\tilde{\zeta}$ in $t\in(0,T(\alpha^{*}))$, hence also the convergence of $\bar{\zeta}(t)$. Hence, Proof of the Lagrangian analyticity of IIE with convergence radius $t\in(0,T(\alpha^{*}))$ is complete. ∎