### v3

# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

Abstract
In this paper, we study the inhomogeneous incompressible Euler equation (IIE in short) from a Lagrangian perspective. We establish a geodesic description of this equation and discuss the associated geometric structures. We also find the derivation of IIE from the Hamilton-Pontryagin action principle and derive the corresponding Lagrangian formulation. A byproduct is a new vorticity formulation of IIE. We also prove the Lagrangian analyticity of IIE using our Lagrangian representation formula.
1 Introduction
The geometry of hydrodynamics has been a vivid research direction since the seminal work of Arnold [2] and Ebin-Marsden [18], where the motion of perfect fluids in a smooth bounded domain $D$ is interpreted as geodesic in the infinite dimensional Riemannian manifold $\mathsf{SDiff}(D)$ of volume-preserving diffeomorphisms of $D$, which is equipped with the right-invariant Riemannian metric given by the kinetic energy of the fluid. This geometric perspective has been extensively generalized since then; we refer to [1][24][30] for further development of this topic, and [26] for a recent survey on geometric hydrodynamics including compressible fluids.
From a mechanical viewpoint, Arnold’s geodesic formulation is a natural infinite-dimensional generalization of the classical least action principle in Lagrangian mechanics. Flipping to the Hamiltonian side, we have an analogous Hamilton-Pontryagin type variational principle for hydrodynamic models (see [6][25]), which leads to the Lagrangian-Eulerian formulation of fluid equations [12]. We remark that the above perspective admits a natural generalization which effectively applies to viscous hydrodynamic models. We refer to [10] for the Navier-Stokes case and to [33] for an extended discussion. However, all the above discussions are restricted to fluid models with mechanical Lagrangians, where velocity and position dependence are decoupled and appear only in kinetic and potential energy respectively.
A natural example which falls outside the class of mechanical Lagrangians is the following inhomogeneous version of the incompressible Euler equation (IIE in short), which is an extensively studied model in past decades:
$\left\{\begin{aligned} &\partial_{t}(\rho u)+\nabla\cdot(\rho u\otimes u)+\nabla p=0\\ &\nabla\cdot u=0\\ &\partial_{t}\rho+u\cdot\nabla\rho=0\end{aligned}\right$. (1)
The above equation describes the motion of incompressible fluids with variable density. Here, the inhomogeneity of the fluid forces the kinetic energy density to depend on both position and velocity variables:
$\frac{1}{2}\lvert u\rvert^{2}\rho=\frac{1}{2}\lvert u\rvert^{2}\rho_{0}\circ X^{-1},\quad\dot{X}=u(X).$
The above equation was systematically studied in the celebrated monograph [28]. Analysis of the equation has been done in various functional settings in a series of recent works, see [14][15][13]. The main focus of the works listed above is to address the well-posedness issues in various function spaces and to analyze IIE from an Eulerian viewpoint using Harmonic analysis tools. However, contrary to the well-known Yudovich theorem of 22-$d$ homogeneous Euler’s equation, the global well-posedness issue of IIE in 22d is still open. In recent years, there have also been works devoted to Beale–Kato–Majda-type (cf.[5] ) blow-up/continuation criteria for IIE, we refer the readers to [3][9][21].
However, different from the analytical side, the geometric structure and Lagrangian perspective of the inhomogeneous Euler equation is much less studied. To the best of the author’s knowledge, the only relevant literature in this direction is [32], which studies the geodesic formulation of IIE on $\mathsf{SDiff}(D)$, as a generalization of the celebrated geodesic formulation of the incompressible Euler equation. The Hamiltonian aspects and Lagrangian formulation (a Weber-type formula, see [12]) of IIE, to the best of the author’s knowledge, are absent in the current literature.
To fill the gap in the literature, in this paper we establish a Lagrangian approach to the analysis of IIE, which stems from geometric consideration. We discuss the Hamiltonian formulation of the equation and derive a Lagrangian-Eulerian representation of the solution of IIE. The advantages of such a Lagrangian perspective are twofold: First, the Lagrangian formulation motivates us to find a vorticity formulation of IIE, which totally eliminates the pressure; Second, benefiting from the Lagrangian perspective, we obtain an explicit recursive formula for the time-series coefficients of the Lagrangian displacement, and demonstrate Lagrangian analyticity of the solution accordingly.
We believe that the vorticity formulation we obtained for IIE may uncover new approaches to the study of IIE. The key takeaway here is that: Geometrically, vorticity in incompressible fluid models is naturally given by the curl of the momentum rather than the velocity, due to the dual Lie transport (coadjoint motion) nature of momentum, together with the following commutation relation, which encodes the vorticity transport:
$\nabla\times\mathscr{L}_{u}^{*}=\mathscr{L}_{u}\nabla\times,\quad\mathscr{L}_{u}^{*}v:=u\cdot\nabla v+\nabla^{*}uv.$
Thus, there is a natural difference between the vorticity of homogeneous and inhomogeneous quantities, based on geometric considerations. As the vorticity formulation plays an important role in the study of various problems of incompressible Euler equation, for instance the finite time blow up, continuation criteria and stability of steady states, the author believes that our vorticity formulation may be helpful in investigating analogous problems for IIE. In our companion paper [35], based on our vorticity formulation, we study the lifespan and continuation criteria of IIE via a Lagrangian approach.
The structure of this paper is organized as follows: In section 2, we study the geodesic formulation of IIE with respect to a weighted Riemannian metric on $\mathsf{SDiff}(D)$, where we interpret the weighted Leray projector as the second-fundamental form. The sectional curvature and Lagrangian stability are also discussed. In section 3, we derive IIE from a Hamilton-Pontryagin type variational principle, and obtain a Lagrangian-Eulerian formulation of IIE as a Lagrangian formula for the solution of the Hamiltonian ODE given by the variational principle. We also derive the vorticity formulation from our geometric construction, and recover the local existence results in our framework. Finally, in section 4, using our Lagrangian representation, we prove the time-analyticity of Lagrangian trajectories of the IIE.
Acknowledgements
The author thanks V.Vicol for stimulating discussions which initiate the study of Lagrangian analyticity of IIE. The author thanks T.Drivas for pointing out the ongoing work [16] and for helpful discussions. The author thanks B.Khesin, H.N.Lopes, M.Lopes, A.Mazzucato and A.Novikov for stimulating discussions and valuable suggestions. This work was partially supported by the US National Science Foundation grants DMS-1909103 and DMS-2206453.
2 Geometry of $\mathsf{SDiff}(D)$ with Weighted Metric
In this section, we study the Riemannian structure of the volume-preserving diffeomorphism group, which is equipped with a weighted Riemannian metric and naturally fits the inhomogeneous incompressible Euler equation. Our presentation in this section mainly follows and generalizes the discussions in the unpublished notes [17], which generalizes constrained geodesic, second fundamental form and sectional curvature of level set submanifold from finite to infinite dimension.
2.1 Some Conventions
We fix some frequently used notations. Let $D$ be the physical domain of our fluids. Unless otherwise specified, we assume $D=\mathbb{T}^{d}$, that is the unit box $0,1]^{d}$ equipped with periodic boundary condition. We remark that most our results also hold in the setting of smooth compact Riemannian manifold or $\mathbb{R}^{d}$, and works under $H^{s+1}$ Sobolev regularity with $s>d/2$, after suitable technical modification (e.g. replacing mean-zero condition with slip boundary condition for compact manifold $D$, or with decay condition for $D=\mathbb{R}^{d}$).
Throughout this paper, we denote by $\mathsf{Diff}(D)$($\mathsf{SDiff}(D)$) the Lie group of (volume-preserving) diffeomorphisms in $D$. We denote by $\mathfrak{X}(D)$, $\mathfrak{X}_{0}(D)$, $\mathfrak{X}_{\sigma}(D)$ the space of smooth vector fields, mean-free vector fields and divergence-free vector fields on $D$.
In this section, we will focus on the geometric formulation of IIE and only consider the smooth case on torus for simplicity. Later, for analytical results like well-posedness and analyticity, we will work in Sobolev regularity class. We will use the notation $H_{0}^{k}(D)$ for space of mean-zero $H^{k}(D)$ functions, and we will frequently use the notation $H^{s}$ without specifying the domain and range of the function. We denote by $\mathbf{P}$ the standard Leray projector on $\mathfrak{X}(D)$, which projects vector fields to its div-free part. In compact form:
$\mathbf{P}v=v-\nabla\Delta^{-1}\nabla\cdot v.$
It is well-known that $\mathbf{P}$ is a zero-order multiplier that preserves the mean value, Hence, it is bounded on each Sobolev space $H_{0}^{k}(D)$ and $\mathbf{P}:\mathfrak{X}_{0}(D)\to\mathfrak{X}_{0}(D)\cap\mathfrak{X}_{\sigma}(D)$.
We will use the letter $X$ for elements in $\mathsf{SDiff}(D)$, and letter $A$ for the inverse of $X$, which is the back-to-label map associated with flow map $X$ in fluid mechanics context.
We denote by $\mathsf{Dens}(D)$ the space of smooth densities in $D$, that is:
$\mathsf{Dens}(D)=\bigg\{\rho\in C^{\infty}(D):\int_{D}\rho(x)dx=1\bigg\}.$
We will fix a reference density $\rho_{0}\in\mathsf{Dens}(D)$, which is required to satisfy the following assumption throughout this paper:
Assumption 2.1 (Assumption (A)).
There exists $C_{1}>1$ such that
$0<C_{1}^{-1}\leq\rho_{0}(a)\leq C_{1}<\infty\quad\text{ for all }a\in D.$
The above assumption 2.1 is posed to avoid concentration or vacuum in fluid density. We denote by
$\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A,\quad X\in\mathsf{SDiff}(D),$
which is the dynamical density transported by some rearrangement $X$. We notice that the assumption 2.1 is propagated under the action of rearrangements. Hence, also propagated by transport equation. We may sometimes add the subscript $X_{t}$ to highlight its time dependence.
We will frequently deal with variable coefficient second-order elliptic operators. Fix a $\rho\in\mathsf{Dens}(D)$, we consider the following weighted Laplacian operator and the associated Poisson equation:
$\Delta_{\rho}\varphi:=\nabla\cdot(\rho^{-1}\nabla\varphi)=f\quad\text{ in }D.$
Assumption 2.1 guaranties that $\Delta_{\rho}$ is an uniformly elliptic operator. Here, we assume the forcing term $f$ satisfies the mean-zero condition:
$\int_{D}f(x)dx=0.$
Then, $\Delta_{\rho}$ is invertible in the space of mean-zero smooth functions. We will denote by $\mathcal{G}_{\rho}$ the Green’s function associated with $\Delta_{\rho}$, and the following holds:
$\int_{D}\mathcal{G}_{\rho}(x,y)dx=0,\quad\mathcal{G}_{\rho}(x,y)=\mathcal{G}_{\rho}(y,x),\quad\varphi(x)=\int_{D}\mathcal{G}_{\rho}(x,y)f(y)dy.$
We will use the letter $C$ to denote positive constants which may depend on additional parameters and change from line to line. $A\lesssim B$ means there exists $C>0$ which might depend on other harmless parameters such that $A\leq CB$.
We end this subsection with the well-known Jacobi-Liouville formula, which will be repeatedly used in the sequel.
Lemma 2.2.
Let $\mathbf{M}:[0,1]\to\mathsf{GL}(d)$ be a smooth curve in the general linear group $\mathsf{GL}(d)$, then we have:
$\frac{d}{dt}\det(\mathbf{M})=\mathsf{Tr}(\mathbf{M}^{-1}\dot{\mathbf{M}})\det(\mathbf{M});$ (2)
and
$\frac{d}{dt}\mathbf{M}^{-1}=-\mathbf{M}^{-1}\dot{\mathbf{M}}\mathbf{M}^{-1}.$ (3)
Proof of Lemma 2.2.
We first notice (3) is a simple consequence of the chain rule:
$0=\frac{d}{dt}(\mathbf{M}\mathbf{M}^{-1})=\dot{\mathbf{M}}\mathbf{M}^{-1}+\mathbf{M}\frac{d}{dt}\mathbf{M}^{-1}.$
Rearranging the above equality, we obtain (3).
To show (2), the standard approach is to use minor expansion. Here, we provide a simple proof suggested by [27].
Notice that $\det\mathbf{M}$ is the Wronskian of matrix $\mathbf{M}$, since the matrix ODE that $\mathbf{M}$ satisfies reads
$\dot{\mathbf{M}}=\mathbf{M}(\mathbf{M}^{-1}\dot{\mathbf{M}})=:\mathbf{M}\tilde{\mathbf{M}}$,
we apply the Abel’s formula for Wronskian (c.f. [37]) to conclude:
$\frac{d}{dt}\det\mathbf{M}=\mathsf{Tr}(\tilde{\mathbf{M}})\det\mathbf{M}_{t}=\mathsf{Tr}(\mathbf{M}^{-1}\dot{\mathbf{M}})\det\mathbf{M}.$
2.2 The Riemannian Structure
Our geometric description relies on the interpretation of $\mathsf{SDiff}(D)$ as a level set submanifold of the full diffeomorphism group $\mathsf{Diff}(D)$ with infinite dimension and co-dimension, and the inhomogeneous Euler equation will be derived as geodesic on the submanifold, with respect to some weighted metric. This formalism closely follows [17]. (See also the ongoing work [16], dealing with dynamics on rapidly oscillating constraint submanifold in infinite dimensions.)
More precisely, we rewrite $\mathsf{SDiff}(D)$ as follows:
$\mathsf{SDiff}(D)=\bigcap_{a\in D}\{X\in\mathsf{Diff}(D):f_{a}(X)=0\},\quad f_{a}(X)=\langle\det(\nabla X)-1,\delta_{a}\rangle_{*}.$
Here, the duality pairing $\langle\cdot,\cdot\rangle_{*}$ is the Riesz pairing between $C_{0}$ functions and measures.
Now for $X\in\mathsf{SDiff}(D)$, to see the tangential direction of the level sets of $f_{a}$, we apply (2) to compute
$\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}f_{a}(X+\varepsilon\delta X)=\langle\mathsf{Tr}((\nabla X)^{-1}\nabla\delta X)\det(\nabla X),\delta_{a}\rangle_{*}$
$\nabla\cdot(\delta X\circ X^{-1})](X(a))=0,$
which implies that
$T_{X}\mathsf{SDiff}(D)=\{v\circ X:v\in\mathfrak{X}_{\sigma}(D)\}.$
Define the following weighted metric:
$\mathsf{G}_{X}(\xi,\eta)=\int_{D}(\xi,\eta)(x)\rho_{0}(x)dx=\int_{D}(\xi\circ X^{-1},\eta\circ X^{-1})\rho(x)dx.$ (4)
Under the above metric, the normal space $N_{X}\mathsf{SDiff}(D)=T_{X}^{\perp}\mathsf{SDiff}(D)$ is given by
$N_{X}\mathsf{SDiff}(D)=\{\rho_{0}^{-1}\nabla\varphi(X):\varphi\in H^{1}(D)\}.$ (5)
Hence, we can define the following map that generates normal vectors:
$\mathcal{N}_{X}:\varphi\in H^{1}(D)\to-\rho_{0}^{-1}\nabla\varphi(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(\varphi\circ X)\in N_{X}\mathsf{SDiff}(D).$ (6)
We now extend the domain of the above normal generator $\mathcal{N}$ to distributions. Let $\mathcal{M}$ be the set of Borel measures on $D$ with finite variation, for any $\mu\in\mathcal{M}$ we define the following co-normal functional
$\mathcal{N}_{X}^{*}(\mu)\in T_{X}^{*}\mathsf{SDiff}(D),\quad\mathcal{N}_{X}^{*}(\mu):=\delta X\in T_{X}\mathsf{Diff}(D)\to\langle\mu,\nabla\cdot(\delta X\circ A)\rangle_{*}.$
which in distributional sense reads:
$\mathcal{N}_{X}^{*}(\mu)=-\nabla\mu(X).$
Moreover, the annihilating property $T_{X}\mathsf{SDiff}(D)\subset\mathsf{Ker}(\mathcal{N}_{X}^{*}(\mu))$ is satisfied.
Moreover, in this context, the musical isomorphism (inertia operator in the Language of mechanics) is $\alpha\to\rho_{0}^{-1}\alpha$. Hence, the vector representative $\mathcal{N}_{X}$ of covector $\mathcal{N}_{X}^{*}$ is:
$\mathcal{N}_{X}(\mu):=\rho_{0}^{-1}\mathcal{N}_{X}^{*}(\mu)=-\rho_{0}^{-1}\nabla\mu(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla(X_{\sharp}\mu),$
which should be interpreted in the distributional sense. Here, $X_{\sharp}\mu$ is the push-forward of measure $\mu$ by diffeomorphism $X$.
In the sequel, we will fix our notation that $A$ will always denote the spatial inverse of diffeomorphism $X$. The following proposition calculates the gradient, Hessian and Gram operator of constraint functions $f_{a}$ in this setting.
Proposition 2.3.
The following statements hold:
  • (I) For any $a\in D$, the gradient $\mathsf{grad}f_{a}$ of $f_{a}$ is:
    $\mathsf{grad}f_{a}(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla\delta_{a}=\mathcal{N}_{X}(\delta_{X(a)})=\mathcal{N}_{X}(\delta_{a}\circ A)$ (7)
  • (II) For any $a\in D$, the Hessian $\mathsf{Hess}f_{a}$ of $f_{a}$ is
    $\mathsf{Hess}f_{a}(X)(\xi,\eta)=-\mathsf{Tr}[\nabla(\xi\circ A)\nabla(\eta\circ A)](X(a)),\quad\forall\xi,\eta\in T_{X}\mathsf{SDiff}(D)$ (8)
  • (III) For any $a,a^{\prime}\in D$, the Gram operator $(Jf)_{a,a^{\prime}}$ is
    $(Jf)_{a,a^{\prime}}:=\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\mathsf{grad}f_{a^{\prime}}(X))=-\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(a)).$ (9)
    While the inverse $(Jf)_{a,a^{\prime}}^{-1}$ reads
    $(Jf)_{a,a^{\prime}}^{-1}(X)=\mathcal{G}_{\rho}(X(a),X(a^{\prime})).$ (10)
    where $\Delta_{\rho}=\nabla\cdot(\rho^{-1}\nabla)$ and $\mathcal{G}_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$, under mean-zero condition.
Proof of Proposition 2.3.
The proof follows from straightforward computation. For (I) we have
$\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\delta X)=\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}f_{a}(X_{\varepsilon})=\langle\nabla\cdot(\delta X\circ A)\circ X,\delta_{a}\rangle_{*}.$
$\langle\delta X\circ A,\nabla(\delta_{a}\circ A)\rangle_{*}=-\langle\delta X,(\nabla^{*}X)^{-1}\nabla\delta_{a}\rangle_{*}$
So we have for any $\delta X\in T_{X}\mathsf{SDiff}(D)$:
$\int_{D}(\mathsf{grad}f_{a}(X),\delta X)\rho_{0}(x)dx=-\int_{D}\delta X\cdot(\nabla^{*}X)^{-1}\nabla\delta_{a}dx.$
Hence, we conclude
$\mathsf{grad}f_{a}(X)=-\rho_{0}^{-1}(\nabla^{*}X)^{-1}\nabla\delta_{a}$ (11)
which completes the proof of (I). For (II), picking any vector fields $u,v\in\mathfrak{X}_{\sigma}$ and denoting $X_{\varepsilon}:=X+\varepsilon v\circ X$, we have:
$\mathsf{Hess}f_{a}(X)(u\circ X,v\circ X)=\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}\mathsf{G}_{X}(\mathsf{grad}f_{a}(X_{\varepsilon}),u\circ X).$
$\int_{D}-\rho_{0}^{-1}\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}(\nabla^{*}X_{\varepsilon})^{-1}\nabla\delta_{a}\cdot u(X)\rho_{0}(x)dx$
$\int_{D}(\nabla^{*}X)^{-1}\nabla^{*}(v\circ X)(\nabla^{*}X)^{-1}\nabla\delta_{a}\cdot u(X)dx.$
$\int_{D}\nabla\delta_{a}(\nabla Au\cdot\nabla v)(X)dx$
$\int_{D}\delta_{a}\nabla\cdot[A_{\sharp}(u\cdot\nabla v)]dx$
$\nabla\cdot(u\cdot\nabla v)X(a)=-\mathsf{Tr}(\nabla u\cdot\nabla v)(X(a)).$
which yields (II). Now for (III), we compute:
$\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\mathsf{grad}f_{a^{\prime}}(X))=\int_{D}(\nabla^{*}X)^{-1}\nabla\delta_{a}(x)(\nabla^{*}X)^{-1}\nabla\delta_{a^{\prime}}(x)\rho_{0}^{-1}(x)dx$
$\int_{D}\nabla(\delta_{a}\circ A)(X)\nabla(\delta_{a^{\prime}}\circ A)(X)\rho_{0}^{-1}(x)dx$
$\int_{D}\nabla(\delta_{a}\circ A)\nabla(\delta_{a^{\prime}}\circ A)\rho_{0}^{-1}\circ Adx$
$\int_{D}(\delta_{a}\circ A)\nabla\cdot[\rho^{-1}\nabla(\delta_{a^{\prime}}\circ A)]$
$\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(a)).$
Meanwhile, the inverse $(Jf)_{a,a^{\prime}}^{-1}$ satisfies:
$\int_{D}(Jf)_{a,z}^{-1}(X)(Jf)_{z,a^{\prime}}(X)dz=\delta(a-a^{\prime}),$
which implies
$\delta(a-a^{\prime})=-\int_{D}(Jf)_{a,z}^{-1}(X)\Delta_{\rho}(\delta_{a^{\prime}}\circ A)(X(z))dz$
$\Delta_{\rho}[(Jf)_{a,A(z)}^{-1}(X)](X(a^{\prime})).$
Hence, we conclude:
$\Delta_{\rho}[(Jf)_{a,A(z)}^{-1}(X)](z)=\delta(X(a)-z).$
As a consequence, we have
$(Jf)_{a,a^{\prime}}^{-1}(X)=-\mathcal{G}_{\rho}(X(a),X(a^{\prime})),$
where $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$. ∎
2.3 Geodesic Equation
We prove in this section the following theorem:
Theorem 2.4.
Let $X$ be a geodesic on infinite dimensional Riemannian manifold $(\mathsf{SDiff}(D),\mathsf{G})$ with $X(0)=\mathbf{Id}$. Then, the following statements hold:
  • (I)
    There exists a pressure field $p$ such that
    $\rho_{0}\ddot{X}=-\nabla p\circ X.$ (12)
  • (II)
    Let $u$ be the velocity field of $X$, then
    $p=-\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u),\quad\rho=X_{\sharp}\rho_{0}.$ (13)
  • (III)
    For $u=\dot{X}\circ X^{-1}$ and $\rho=X_{\sharp}\rho_{0},$ $(u,\rho,p)$ solves the IIE (1).
Proof of Theorem 2.4.
  • (I)
    By D’Alembert principle, we have $\ddot{X}\in T_{X}^{\perp}\mathsf{SDiff}(D)$. Hence, there exists a scalar function $\lambda:[0,T]\times D\to\mathbb{R}$ such that for any $x\in D$:
    $\rho_{0}(x)\ddot{X}(x)=\int_{D}\lambda(t,z)(\nabla^{*}X)^{-1}(x)\nabla_{z}\delta_{x}(z)dz$ (14)
    $(\nabla^{*}X)^{-1}(x)\int_{D}\underbrace{\nabla_{z}(\delta_{z}(x)\lambda(t,z))}_{=0\quad a.e.}-\delta_{z}(x)\nabla\lambda(t,z)dz$
    $(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(z)\nabla(\underbrace{\lambda\circ A}_{:=p})(X(z))\delta_{z}(x)dz$
    $(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(z)\nabla p(X(z))\delta_{x}(z)dz=-\nabla p(t,X),$
  • (II)
    We take a closer look at the Lagrange multiplier $\lambda$. Since for any $a\in D$ we have:
    $\mathsf{G}_{X}(\mathsf{grad}f_{a}(X),\dot{X})=0.$ (15)
    Differentiating (15) in time and denoting by $u=\dot{X}_{t}\circ X_{t}^{-1}$, we have for any fixed $a\in D$:
    $0=-\frac{d}{dt}\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\dot{X}_{t}(x)dx.$
    $\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\ddot{X}_{t}(x)dx+\int_{D}(\nabla^{*}X_{t})^{-1}\nabla^{*}(\dot{X}_{t}(x))(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}\cdot\dot{X}_{t}(x)dx$
    $\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\ddot{X}_{t}(x)dx+\int_{D}\nabla\delta_{a}(x)(A_{t})_{\sharp}(u\cdot\nabla u)dx$
    $\int_{D}(\nabla^{*}X_{t})^{-1}\nabla\delta_{a}(x)\cdot\ddot{X}_{t}(x)dx-\int_{D}\delta_{a}(x)[\nabla\cdot(u\cdot\nabla u)](X_{t}(x))dx.$
    Plugging (14) into the above equality, we have the following.
    $\nabla\cdot(u\cdot\nabla u)](X_{t}(a))=-\int_{D}(\nabla^{*}X_{t})^{-1}(x)\nabla\delta_{a}(x)\int_{D}\rho_{0}^{-1}(x)\lambda(z)(\nabla^{*}X_{t})^{-1}(x)\nabla_{z}\delta_{x}(z)dzdx$
    $\int_{D}\int_{D}\nabla(\delta_{a}\circ A_{t})(X_{t}(x))\rho_{0}^{-1}(x)\lambda(z)(\nabla^{*}X_{t})^{-1}(x)\nabla_{z}\delta_{x}(z)dzdx$
    $\int_{D}\int_{D}\nabla(\delta_{a}\circ A_{t})\rho_{t}^{-1}(x)\nabla^{*}A_{t}(x)\nabla\lambda(z)\delta_{A_{t}(x)}(z)dzdx$
    $\int_{D}\nabla(\delta_{a}\circ A_{t})\rho_{t}^{-1}(x)\nabla(\lambda\circ A_{t})(x)dx$
    $\nabla\cdot(\rho_{t}^{-1}\nabla p)(X_{t}(a)),\quad p:=\lambda\circ A_{t}.$
    Hence, we conclude:
    $p=-\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u),\quad\lambda=p\circ X.$ (16)
  • (III)
    The above geodesic equation (12) in Eulerian coordinates reads:
    $\rho(\partial_{t}u+u\cdot\nabla u)-\nabla\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla u)=0$
    Together with $\nabla\cdot u=0$ and $\rho$ transported by $u$, we readily verify that IIE (1) is the geodesic equation on $\mathsf{SDiff}(D)$ with our weighted metric $\mathsf{G}$.
2.4 Second Fundamental Form and Curvature Tensor
Viewing $\mathsf{SDiff}(D)$ as a submanifold of $\mathsf{Diff}(D)$ of infinite dimension and codimension, it makes sense to talk about the second fundamental form of $\mathsf{SDiff}(D)$, which admits a concrete description as the projection of the convective acceleration onto the normal direction. To see this, we define for any given $\rho\in\mathsf{Dens}(D)$ the following operator
$\mathbf{Q}_{\rho}:w\to\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot w$ (17)
The next lemma shows that $\mathbf{Q}_{\rho}$ is an orthogonal projector in weighted $L^{2}$-sense.
Proposition 2.5.
Let $\rho\in\mathsf{Dens}(D)$ and smooth vector field $w\in\mathfrak{X}(D)$, then the following statements hold:
  • (I)
    There exists a unique div-free vector field $v$ and $p\in C^{\infty}(D)$ such that
    $w=\rho v+\nabla p.$ (18)
    moreover, $v=(\mathbf{Id}-\mathbf{Q}_{\rho})(\rho^{-1}w)=:\mathbf{P}_{\rho}(\rho^{-1}w)$. Here, $v$ is not necessarily mean-zero, but $\rho v\in\mathfrak{X}_{0}(D)$ iff $w\in\mathfrak{X}_{0}(D)$.
  • (II)
    Let $\mathcal{H}_{\rho}:=\{\rho^{-1}\nabla\varphi:\varphi\in C^{\infty}(D)\}$, then we have $\mathsf{Ker}(\mathbf{Id}-\mathbf{Q}_{\rho})=\mathcal{H}_{\rho}$. In particular, we have $\mathcal{H}_{\rho}\perp\mathsf{Ran}(\mathbf{P}_{\rho})$ in the sense of $L_{\rho}^{2}$.
Proof of Proposition 2.5.
  • (I)
    Given vector field $w$, for any given $\rho\in\mathsf{Dens}(D)$ we write
    $w=\rho\underbrace{[\rho^{-1}w-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)]}_{:={\mathbf{P}}_{\rho}(\rho^{-1}w).}+\underbrace{\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)}_{=\rho\mathbf{Q}_{\rho}(\rho^{-1}w)}$ (19)
    Let $v=\mathbf{P}_{\rho}(\rho^{-1}w)$ and $p=\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}w)$, we find the decomposition. Moreover, assume $w=\rho v^{\prime}+\nabla p^{\prime}$ for some div-free $v^{\prime}$ and $p^{\prime}\in C^{\infty}(D)$, then $p^{\prime}$ can be recovered from solving the following elliptic problem:
    $\Delta_{\rho}p^{\prime}=\nabla\cdot(\rho^{-1}\nabla p^{\prime})=\nabla\cdot(\rho^{-1}w)-\underbrace{\nabla\cdot v^{\prime}}_{=0}=\nabla\cdot(\rho^{-1}w).$
    Uniform ellipticity of $\Delta_{\rho}$ implies the uniqueness of solution $p^{\prime}$, hence the uniqueness of $v^{\prime}$.
  • (II)
    Picking any $\varphi\in C^{\infty}(D)$, we compute $\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)$ as follows:
    $\mathbf{P}_{\rho}(\rho^{-1}\nabla\varphi)=\rho^{-1}\nabla\varphi-\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot(\rho^{-1}\nabla\varphi)=0,$
    which guaranties $\mathcal{H}_{\rho}\subset\mathsf{Ker}(\mathbf{P}_{\rho})$. Conversely, assume we have $\mathbf{P}_{\rho}\xi=0$, then by the decomposition:
    $\xi=\mathbf{Q}_{\rho}\xi=\rho^{-1}\nabla\Delta_{\rho}^{-1}\nabla\cdot\xi\in\mathcal{H}_{\rho}.$
    which completes the proof of (II).
The above proposition gives a dynamical decomposition of vector fields into vertical ($\mathfrak{X}_{\sigma}$) and horizontal ($\mathcal{H}_{\rho}$) directions, from the ambient $\mathsf{Diff}(D)$ viewpoint.
We now connect the second fundamental form $\Pi$ of $\mathsf{SDiff}(D)$ w.r.t metric $\mathsf{G}$ and the above weighted projector $\mathbf{Q}_{\rho}$:
Theorem 2.6.
The second fundamental form $\Pi_{X}$ of $\mathsf{SDiff}(D)$ at $X$ is given by
$\Pi_{X}(u\circ X,v\circ X)=\mathbf{Q}_{\rho}(u\cdot\nabla v),\quad\forall u,v\in\mathfrak{X}_{\sigma}(D)\text{ and }X\in\mathsf{SDiff}(D).$ (20)
Proof of Theorem 2.6.
Formula (20) follows from straightforward computation. Mimicking the definition for second fundamental form in finite dimension:
$\Pi_{X}(u\circ X,v\circ X)(x)=\int_{D\times D}(Jf)_{y,z}^{-1}(X)\mathsf{Hess}f_{z}(X)(u\circ X,v\circ X)\mathsf{grad}f_{y}(X(x))dydz$
$\int_{D}[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(y))\rho_{0}^{-1}(x)(\nabla^{*}X(x))^{-1}\nabla_{y}\delta_{x}(y)dy$
$\rho_{0}^{-1}(x)(\nabla^{*}X)^{-1}(x)\int_{D}[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(y))\nabla_{y}\delta_{x}(y)dy$
$\rho_{0}^{-1}(x)(\nabla^{*}X)^{-1}(x)\int_{D}\nabla^{*}X(y)\nabla[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(y))\delta_{x}(y)dy$
$\rho_{0}^{-1}(x)\nabla[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)](X(x))$
$(\rho^{-1}\nabla[\Delta_{\rho}^{-1}\nabla\cdot(u\cdot\nabla v)])(X(x))=\mathbf{Q}_{\rho}(u\cdot\nabla v)(X).$
Remark 2.7.
$\Pi_{X}$ is symmetric on $T_{X}\mathsf{SDiff}(D)\times T_{X}\mathsf{SDiff}(D)$, since
$\Pi_{X}(u\circ X,v\circ X)-\Pi_{X}(v\circ X,u\circ X)=\mathbf{Q}_{\rho}([u,v])(X)=0.$
Here, we use the fact that $\nabla\cdot[u,v]=0$ provided $u,v\in\mathfrak{X}_{\sigma}(D)$.
With the above explicit formula for $\Pi$, we can now compute the curvature tensor as follows: For tangent vector $\xi=u\circ X,\eta=v\circ X\in T_{X}\mathsf{SDiff}(D)$ with $u,v\in\mathfrak{X}_{\sigma}(D)$, we have
$\mathcal{C}_{\xi,\eta}=\int_{D}[\Pi_{X}(\xi,\xi)(z)\cdot\Pi_{X}(\eta,\eta)(z)-\lvert\Pi_{X}(\xi,\eta)(z)\rvert^{2}]\rho_{0}(z)dz$
$\int_{D}[\mathbf{Q}_{\rho}(u\cdot\nabla u)\mathbf{Q}_{\rho}(v\cdot\nabla v)-\lvert\mathbf{Q}_{\rho}(u\cdot\nabla v)\rvert^{2}]\rho(z)dz.$
In view of the above curvature formula, we recover the Lagrangian instability of shear flows, as steady states of the homogeneous Euler equation.
3 Hamiltonian and Vorticity Formulation
In this section, we flip to the Hamiltonian side of the above geometric picture of IIE and derive the Lagrangian-Eulerian formulation of IIE from Hamilton’s action principle. The Hamiltonian formulation of the motion of ideal fluids reads coadjoint motion of one-form on $\mathfrak{X}_{\sigma}^{*}(D)$, lies in the framework of general Lie-Poisson equations [30], while the Leray projector essentially becomes the musical isomorphism from cotangent to tangent vectors (see [31][1]). Identifying $\mathfrak{X}_{\sigma}(D)$ with the space of scalar(vector, resp.) potentials by $\nabla^{\perp}(\nabla\times$,resp.) in $2(3$-d,resp.), we obtain the vortex dynamics.
Here, we extend the above geometric picture to the inhomogeneous setting. A Hamilton-Pontryagin type variational principle and a companion Lagrangian-Eulerian formulation is established, which generalizes the Weber’s formula for perfect fluids in [12]. We then establish the vorticity formulation for IIE in simply connected domain and torus respectively. The reason why we need to treat torus case separately is because the mean value of $\rho u$ is conserved, instead of $u$, hence a global stream function representation for $u$ is not applicable in torus case. A proof of local existence of strong solutions for IIE is obtained based on our vorticity formulation.
3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
We now lift ourselves to the cotangent bundle of $\mathsf{SDiff}(D)$ where the Hamiltonian equation naturally lives and we will derive it in our setting. To this end, we first write down the kinetic energy as a functional on tangent bundle:
$L:(X,\dot{X})\in T_{X}\mathsf{SDiff}(D)\to L(X,\dot{X}):=\frac{1}{2}\int_{D}\lvert\dot{X}(t,x)\rvert^{2}\rho_{0}(x)dx$ (21)
It is not hard to show that the action-minimizing problem of $L$, augmented by the incompressibility constraint $\det(\nabla X)=1$, is precisely solved by the solution of equation (1). We refer the interested readers to [29] for study of a relaxed geodesic problem in the sense of Brenier’s generalized flow (cf. [7], [8]).
Abusing the notation a little bit, we now write the Lagrangian $L$ in an equivalent form:
$L(u,X)=\frac{1}{2}\int_{D}\lvert u\rvert^{2}(X_{\sharp}\rho_{0})(x)dx$ (22)
Let $D=\mathbb{R}^{d}$ or $\mathbb{T}^{d}$, we now state the following:
Theorem 3.1 ((Hamilton-Pontryagin Principle)).
Fix $T>0$ and $\rho_{0}\in\mathsf{Dens}(D)$ satisfying assumption 2.1, we assume the triplet
$(X,u,\xi)\in C^{1}([0,T];\mathsf{SDiff}(D)\times\mathfrak{X}_{\sigma}(D)\times\mathfrak{X}(D))$
is a critical point of the augmented action
$\Phi(u,X,\xi)=\int_{0}^{T}L(u,X)-\langle\xi,\dot{X}-u(X)\rangle_{L_{\rho_{0}}^{2}}dt.$ (23)
Then the following statements hold:
  • (I)
    $(u_{t},\rho_{0}\circ X_{t}^{-1})$ solves the incompressible inhomogeneous Euler equation (1).
  • (II)
    Consider the following controlled Hamiltonian:
    $H(\xi,X,u)=L(u,X)+\langle\xi,u(X)\rangle_{L_{\rho_{0}}^{2}}.$
    Then $(X,u,\xi)$ solves the associated Hamiltonian ODE:
    $\dot{X}=\frac{\partial H}{\partial\xi},\quad\dot{\xi}=-\frac{\partial H}{\partial X},\quad\frac{\partial H}{\partial u}=0.$ (24)
  • (III)
    Given the initial data $(u_{0},\xi_{0})$ satisfying compatibility condition
    $u_{0}=\mathbf{P}_{\rho_{0}}\xi_{0},\quad\text{equivalently }\xi_{0}=u_{0}+\rho_{0}^{-1}\nabla q_{0}.$
    the solution of (24) is given by the following Lagrangian-Eulerian system:
    $\left\{\begin{aligned} &\dot{X}_{t}=u(t,X_{t}),A_{t}=X_{t}^{-1};\\ &\rho(t,\cdot)=\rho_{0}\circ A_{t};\\ &u_{t}=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg].\end{aligned}\right$. (25)
Proof of Theorem 3.1.
  • (I)
    As a reminder of our notation, we denote by $A=X^{-1}$ and $\rho:=\rho_{0}\circ A$. We compute the first variation of $\Phi$. We consider smooth perturbations $\delta u\in C^{1}([0,T];\mathfrak{X}_{\sigma}(D))$, $\delta X\in C_{c}^{1}([0,T];\mathsf{SDiff}(D))$ and $\delta\xi\in C^{1}([0,T];\mathfrak{X}(D))$ of $\Phi$, respectively, and compute the corresponding variational equations as follows:
    • (i)
      The variation in the variable $u$ is given by
      $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u+\varepsilon\delta u,X,\xi)=\int_{0}^{T}\int_{D}\langle u\rho,\delta u\rangle-\langle\xi,\delta u(X)\rangle_{L_{\rho_{0}}^{2}}dxdt.$ (26)
      Hence, vanishing of (26) implies that $\rho(u-\xi\circ A)$ is a gradient, i.e. there exists some scalar function $q$ such that
      $u\rho=(\rho_{0}\xi)\circ A+\nabla q.$ (27)
    • (ii)
      The variation in variable $X$ is given by
      $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u,X+\varepsilon\delta X,\xi)$
      $\int_{0}^{T}\int_{D}\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\frac{1}{2}\lvert u\rvert^{2}\rho_{0}\circ X_{\varepsilon}^{-1}+\langle\xi,\dot{X}+\varepsilon\dot{\delta X}\rangle_{L_{\rho_{0}}^{2}}-\langle\xi,u(X+\varepsilon\delta X)\rangle_{L_{\rho_{0}}^{2}}dxdt.$ (28)
      To compute $\frac{d}{d\varepsilon}X_{\varepsilon}^{-1}$, notice:
      $0=\frac{d}{d\varepsilon}(X_{\varepsilon}\circ X_{\varepsilon}^{-1})=\delta X\circ X_{\varepsilon}^{-1}+\nabla X_{\varepsilon}\circ X_{\varepsilon}^{-1}\cdot\frac{d}{d\varepsilon}X_{\varepsilon}^{-1}.$
      Hence, we have:
      $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}X_{\varepsilon}^{-1}=-[(\nabla X_{\varepsilon})^{-1}\delta X]\circ X_{\varepsilon}^{-1}\bigg\rvert_{\varepsilon=0}=-[(\nabla X)^{-1}\delta X]\circ A.$ (29)
      Therefore, combining (28) and (29) we conclude
      $\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\Phi(u,X+\varepsilon\delta X,\xi)$
      $\int_{0}^{T}\int_{D}-\frac{1}{2}\lvert u\rvert^{2}\nabla\rho_{0}\circ A\cdot[(\nabla X)^{-1}\delta X]\circ A+\langle\xi,\dot{\delta X}\rangle_{L_{\rho_{0}}^{2}}-\langle\xi,\nabla u(X)\delta X\rangle_{L_{\rho_{0}}^{2}}dxdt$
      $\int_{0}^{T}\int_{D}\langle\rho_{0}\dot{\xi}+\rho_{0}\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X,\delta X\rangle dxdt.$ (30)
      By arbitrariness of $\delta X$ in (30), the variational equation in $X$ reads
      $\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0.$ (31)
    • (iii)
      The variation in variable $\xi$ is simply given by
      $\dot{X}=u(X).$ (32)
      Hence, the critical point equation is obtained by putting together (27)(30)(32):
      $\left\{\begin{aligned} &\dot{X}=u(X),\quad A=X^{-1};\\ &\rho_{0}[\dot{\xi}+\nabla^{*}u(X)\xi]+\frac{1}{2}\lvert u\circ X\rvert^{2}\nabla\rho\circ X=0;\\ &u\rho=(\rho_{0}\xi)\circ A+\nabla q,\quad\nabla\cdot u=0;\\ &\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A.\end{aligned}\right$. (33)
    Now denote by $\eta=(\rho_{0}\xi)\circ A$, by method of characteristics, (31) implies
    $\partial_{t}\eta=(\rho_{0}\dot{\xi})\circ A+\nabla(\rho_{0}\xi)\circ A\cdot\partial_{t}A=-\nabla^{*}u(\rho_{0}\xi)\circ A-\frac{1}{2}\lvert u\rvert^{2}\nabla\rho-u\cdot\nabla[(\rho_{0}\xi)\circ A].$
    hence $\eta$ solves the PDE:
    $\partial_{t}\eta+u\cdot\nabla\eta+\nabla^{*}u\cdot\eta+\frac{1}{2}\lvert u\rvert^{2}\nabla\rho=0.$ (34)
    Now, since $u\rho=\nabla q+\eta$, plugging in the equation (34) we obtain
    $\partial_{t}(u\rho-\nabla q)+u\cdot\nabla(u\rho-\nabla q)+\nabla^{*}u\cdot(u\rho-\nabla q)+\frac{1}{2}\lvert u\rvert^{2}\nabla\rho=0.$
    Simple rearrangement yields
    $\partial_{t}(\rho u)+u\cdot\nabla(\rho u)+\nabla\bigg(\frac{1}{2}\lvert u\rvert^{2}\rho-u\cdot\nabla q-\partial_{t}q\bigg)=0.$
    Hence, we let
    $p=\frac{1}{2}\lvert u\rvert^{2}\rho-u\cdot\nabla q-\partial_{t}q$
    and conclude that (33) is equivalent to the system (1) with pressure $p$.
  • (II)
    We compute the partial derivatives $\partial_{\xi}H$, $\partial_{X}H$. Recall the definition of variational derivative:
    $\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi+\varepsilon\delta\xi,X,u)=\int_{D}[\frac{\partial H}{\partial\xi}(\xi,X,u)\cdot\delta\xi]\rho_{0}(a)da,$
    and
    $\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi,X+\varepsilon\delta X,u)=\mathsf{G}_{X}\big(\frac{\partial H}{\partial X},\delta X\big)=\int_{D}[\frac{\partial H}{\partial X}(\xi,X,u)\cdot\delta X]\rho_{0}(a)da;$
    which implies the following Hamilton’s ODE
    $\frac{\partial H}{\partial\xi}=u(X),\quad\frac{\partial H}{\partial X}=\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u(X)\rvert^{2}(\nabla^{*}X)^{-1}\nabla\log\rho_{0};$
    with control
    $\frac{\partial H}{\partial u}=u\rho-(\rho_{0}\xi)\circ A-\nabla q=0.$
    Putting above together, straightforward computation yields that (24) is equivalent to (33).
  • (III)
    We consider the following general version of (31):
    $\dot{w}+\nabla^{*}u(X)w=f(X),\quad w(0,\cdot)=w_{0}.$
    Provided sufficient regularity on $X$, applying Duhamel’s formula we have the following Lagrangian formula for solution $w(t,\cdot)$:
    $w(t,x)=(\nabla^{*}X_{t})^{-1}\bigg(w_{0}(x)+\int_{0}^{t}(\nabla^{*}X_{\tau})f(\tau,X_{\tau}(x))d\tau\bigg)$
    Hence, we obtain the following representation for the solution $\eta$ of (34):
    $\eta(t,x)=\nabla^{*}A_{t}\bigg(\rho_{0}\xi_{0}-\frac{1}{2}\int_{0}^{t}(\nabla^{*}X_{\tau})\lvert u(\tau,X)\rvert^{2}\nabla\rho(\tau,X_{\tau}(x))d\tau\bigg)\circ A_{t}(x)$
    $\nabla^{*}A_{t}\bigg(\rho_{0}\xi_{0}-\frac{1}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}\nabla\rho_{0}(x)d\tau\bigg)\circ A_{t}(x).$ (35)
    We are now ready to conclude:
    $u=\mathbf{P}_{\rho}(\rho^{-1}\eta)=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg].$ (36)
    Combining (32) (36) and $\rho=\rho_{0}\circ A$, we recover the system (25).
3.2 Vorticity Formulation
We now move on to derive a stream-vorticity formulation of (1). We will state a more general functional analytic framework and treat IIE as a specific example. Here we treat the case that $D$ is simply connected and $D=\mathbb{T}^{2}$ separately.
Let $\mathcal{V}$ be a Hilbert space with inner product $\langle\cdot,\cdot\rangle_{\mathcal{V}}$. Let $\mathcal{I}:\mathcal{V}\to\mathcal{V}^{*}$ be the Riesz isomorphism between dual pairing $\langle\cdot,\cdot\rangle_{*}$ of $\mathcal{V},\mathcal{V}^{*}$:
$\langle v,v^{\prime}\rangle_{*}=\langle v,\mathcal{I}^{-1}v^{\prime}\rangle_{\mathcal{V}}$
Let $\mathcal{D}:\mathcal{F}\to\mathcal{V}$ be a linear operator from Hilbert space $\mathcal{F}$ (with dual pairing $\langle,\rangle$) to the Hilbert space $\mathcal{V}$ with closed range, then we have the following pull-back inner product on $\mathcal{F}$:
$\langle\varphi,\psi\rangle_{\mathcal{F}}:=\langle\mathcal{D}\varphi,\mathcal{D}\psi\rangle_{\mathcal{V}}=\langle\mathcal{D}\varphi,\mathcal{I}\mathcal{D}\psi\rangle_{*}=\langle\varphi,\mathcal{D}^{*}\mathcal{I}\mathcal{D}\psi\rangle$ (37)
Denote by $\mathcal{A}:\mathcal{F}\to\mathcal{F}^{*}$ the Riesz isomorphism on $(\mathcal{F},\langle,\rangle_{\mathcal{F}})$ we have for $\psi^{\prime}\in\mathcal{F}^{*}$ with $\mathcal{A}\psi=\psi^{\prime}$:
$\langle\varphi,\mathcal{A}\psi\rangle=\langle\varphi,\mathcal{D}^{*}\mathcal{I}\mathcal{D}\psi\rangle.$ (38)
Hence, we conclude: $\mathcal{A}=\mathcal{D}^{*}\mathcal{I}\mathcal{D}$. While the projection onto the range $\mathsf{Ran}(\mathcal{D})$ is:
$\mathcal{P}=\mathcal{D}\mathcal{A}^{-1}\mathcal{D}^{*}\mathcal{I}:\mathcal{V}\to\mathsf{Ran}(\mathcal{D})$
In our hydrodynamic setting, assume $d=2$ and $D\subset\mathbb{R}^{2}$ is a simply connected smooth domain. We let $\mathcal{V}=L^{2}(D;\mathbb{R}^{d})$ be the space of $L^{2}$ vector fields, $\mathcal{F}=H^{1}(D)/\mathbb{R}$, $\mathcal{F}^{*}=H_{0}^{-1}(D)$ be the space of $H^{-1}$ distributions. Set $\mathcal{D}=\nabla^{\perp}$, hence $\mathcal{D}^{*}=\nabla\times$. The inertia operator in this case is the following uniformly elliptic operator (due to assumption 2.1):
$\mathcal{A}=\nabla\times(\rho\nabla^{\perp})=:\mathbf{L}_{\rho}=\nabla\cdot(\rho\nabla),$
and the projector reads:
$\mathcal{P}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho(\cdot)).$
Here, the inverse $\mathbf{L}_{\rho}^{-1}$ is understood in the sense of Dirichlet boundary condition.
Indeed, decomposing $\rho^{-1}w=v+\rho^{-1}\nabla p$ as in proposition 2.5(I), we have the following identity
$\mathcal{P}(\rho^{-1}w)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho v)=:\mathcal{K}_{\rho}\nabla\times(\rho v).$ (39)
Here, $\mathcal{K}_{\rho}$ is the modified Biot-Savart operator, which plays the role of the right inverse of the $\rho$-weighted curl operator. (39) is a generalization of the following identity in the homogeneous setting:
$\mathbf{P}v=\nabla^{\perp}\Delta^{-1}\nabla\times v,\quad\forall v\in\mathfrak{X}(D).$
Since $d=2$ and $\nabla\cdot v=0$, let $\psi$ be the stream function of $v$, we find:
$\mathcal{P}(\rho^{-1}w)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho\nabla^{\perp}\psi)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\mathbf{L}_{\rho}\psi=\nabla^{\perp}\psi=v.$
Therefore $\mathcal{P}=\mathbf{P}_{\rho}$. Similar calculation shows that $\mathcal{P}=\mathbf{P}_{\rho}$ also holds in $d=3$, if we replace $\nabla^{\perp}$ with $\nabla\times$ and define $\mathbf{L}_{\rho}$ as the following weighted Laplacian for vector fields:
$\mathbf{L}_{\rho}:w\in\mathfrak{X}(D)\to\nabla\times(\rho\nabla\times w).$
We formally unify the above relations in $2$-d in the following commutative diagram.
$\mathcal{F}:=H_{0}^{1}(D)$$\mathcal{V}:=L^{2}_{\rho}(D;\mathbb{R}^{2})$$\mathcal{F}^{*}:=H_{0}^{-1}(D)$$\mathcal{V}^{*}\cong L^{2}_{\rho}(D;\mathbb{R}^{2})$$\scriptstyle{\mathcal{D}=\nabla^{\perp}}$$\scriptstyle{\mathcal{I}:\ v\mapsto\rho v}$$\scriptstyle{\mathbf{P}_{\rho}=\mathcal{D}\mathcal{A}^{-1}\mathcal{D}^{*}\mathcal{I}}$$\scriptstyle{\mathcal{A}^{-1}=\mathbf{L}_{\rho}^{-1}}$$\scriptstyle{\mathcal{K}_{\rho}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}}$$\scriptstyle{\mathcal{D}^{*}=\nabla\times}$ (40)
Following the geometric picture above, we now introduce the vorticity formulation in this simply-connected setting. Let $\eta=\rho u+\nabla q$, we therefore have $\mathbf{P}_{\rho}(\rho^{-1}\eta)=u$. Following our above computation, we have yet another closed-form representation of $\mathbf{P}_{\rho}$:
$\mathbf{P}_{\rho}(\rho^{-1}\eta)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times\eta.$
Now we define $\omega=\nabla\times\eta$, then taking the curl of (24) we obtain the following vorticity equation:
$\partial_{t}\omega+\mathscr{L}_{u}\omega+\nabla\frac{1}{2}\lvert u\rvert^{2}\times\nabla\rho=0.$ (41)
where $\mathscr{L}_{u}$ is the Lie derivative of a scalar/vector field in $2d/3d$ respectively. Hence, in view of the above diagram, we have the following vorticity formulation of (1), where the pressure is deleted:
$\left\{\begin{aligned} &\partial_{t}\omega+\mathscr{L}_{u}\omega+\nabla\frac{1}{2}\lvert u\rvert^{2}\times\nabla\rho=0\\ &u=\mathcal{K}_{\rho}\omega:=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\omega\\ &\partial_{t}\rho+u\cdot\nabla\rho=0\end{aligned}\right$. (42)
We note that (42) is algebraically equivalent to system (25). The boundary condition is $u\cdot\mathbf{n}=0$ on $\partial D$, which is compatible with the Dirichlet boundary condition of the Poisson equation
$\mathbf{L}_{\rho}\psi=\omega,\quad\psi\big\rvert_{\partial D}=0.$ (43)
We now show the conservation of vorticity (Kelvin circulation theorem) in the framework of IIE. In hydrodynamics, conservation laws coming from particle relabeling symmetry (Casimir Invariants in geometry) are known to be the consequence of duality between Lie transport (Lie elements of Lagrangian loops) and dual Lie transport (co-adjoint motion of momentum, cf.[24][33]). Here, for the inhomogeneous fluid, we do not have a full conservation of circulation, due to the forcing term in vortex dynamics. However, the following result holds for the case that $D\subset\mathbb{R}^{2}$ is simply connected:
Theorem 3.2 ((Conservation of Circulation)).
Assume $(u,\rho)$ is a smooth solution of IIE. For any regular value $\alpha>0$ of $\rho_{0}$, assume $E_{\alpha}\subset\{\rho_{0}=\alpha\}$ is a smooth loop, let $D_{\alpha}:=\{\rho_{0}<\alpha\}$ be the smooth bounded domain enclosed by $E_{\alpha}$, then we have
$\oint_{X_{t}(E_{\alpha})}u_{t}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell,$ (44)
which implies
$\int_{X_{t}(D_{\alpha})}\omega_{t}(x)dx=\int_{D_{\alpha}}\omega_{0}(x)dx.$ (45)
Proof of Theorem 3.2.
Thanks to the Lagrangian formula (35), we have for any smooth loop $\gamma$:
$\oint_{X_{t}(\gamma)}\eta_{t}\cdot d\ell=\oint_{\gamma}\nabla^{*}X_{t}\eta_{t}\circ X_{t}\cdot d\ell$
$\oint_{\gamma}\big[u_{0}\rho_{0}+\nabla q_{0}-\nabla\rho_{0}(x)\frac{1}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\big]\cdot d\ell$
$\oint_{\gamma}u_{0}\rho_{0}\cdot d\ell-\frac{1}{2}\int_{0}^{t}\oint_{\gamma}\lvert\dot{X}_{\tau}\rvert^{2}\nabla\rho_{0}\cdot d\ell d\tau.$
Now choose $\gamma=E_{\alpha}$, since $\{\rho_{t}=\alpha\}=X_{t}(E_{\alpha})$, we conclude:
$\oint_{X_{t}(E_{\alpha})}u_{t}\cdot d\ell=\alpha^{-1}\oint_{X_{t}(E_{\alpha})}\eta_{t}\cdot d\ell=\alpha^{-1}\oint_{E_{\alpha}}\rho_{0}u_{0}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell.$
where we use the orthogonality relation: For all smooth $f$ we have
$\oint_{E_{\alpha}}f\nabla\rho_{0}\cdot d\ell=\int f\nabla\rho_{0}\cdot\frac{\nabla^{\perp}\rho_{0}}{\lvert\nabla\rho_{0}\rvert}\mathscr{H}_{E_{\alpha}}^{1}(dx)=0.$
Hence, the proof of (44) is complete. To show (45), simply apply the Stokes theorem to (44) multiplied by $\alpha$.
Remark 3.3 ((Noether Theorem)).
Actually, we may derive the conservation of circulation of IIE as a consequence of Noether’s theorem. By (21), we consider the following action:
$S(\{X_{t}\})=\int_{0}^{T}L(X,\dot{X})dt=\int_{0}^{T}\int_{D}\frac{1}{2}\lvert\dot{X}(t,a)\rvert^{2}\rho_{0}(a)da$
We notice that $S$ is invariant under right action of $\Psi\in\mathsf{Diff}_{\rho_{0}}(D)$, the subgroup of $\mathsf{Diff}(D)$ preserving the initial density $\rho_{0}$. Moreover, by Lemma 2.2, for curve $\Psi_{\varepsilon}\subset\mathsf{Diff}_{\rho_{0}}(D)$ generated by vector field $v$ and starting from identity, we have:
$0=\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\rho_{0}=\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\rho_{0}(\Psi_{\varepsilon})\det(\nabla\Psi_{\varepsilon})=\rho_{0}\nabla\cdot v+v\cdot\nabla\rho_{0}=\nabla\cdot(v\rho_{0})$
Hence, $\rho_{0}$-preserving flows are generated by vector field $v$ such that $\nabla\cdot(v\rho_{0})=0$. Now, combined with the $\mathsf{SDiff}(D)$-constraint of the IIE dynamics, we conclude that $\mathscr{G}=\mathsf{SDiff}(D)\cap\mathsf{Diff}_{\rho_{0}}(D)$ is the right gauge group for the IIE action $S$, which is significantly smaller than $\mathsf{SDiff}(D)$ and fails the full Kelvin circulation theorem. In view of the above computation, we have:
$T_{X}\mathscr{G}=\{v\circ X:\nabla\cdot v=v\cdot\nabla\rho_{0}=0\}.$
Differentiating $S$ with right action of $\{\Psi_{\varepsilon}\}\subset\mathscr{G}$, we compute:
$\frac{d}{d\varepsilon}\bigg|_{\varepsilon=0}S(\{X_{t}\circ\Psi_{\varepsilon}\})=\frac{d}{d\varepsilon}\bigg|_{\varepsilon=0}\int_{0}^{T}\int_{D}\frac{1}{2}\lvert\dot{X}(t,\Psi_{\varepsilon}(a))\rvert^{2}\rho_{0}(a)da$
$\int_{0}^{T}\int_{D}\dot{X}(t,a)[\nabla\dot{X}(t,a)v(a)]\rho_{0}(a)dadt$
$\int_{0}^{T}\int_{D}\frac{d}{dt}[\dot{X}(t,a)(\nabla X(t,a)v)]\rho_{0}(a)dadt-\int_{0}^{T}\int_{D}\rho_{0}(a)\ddot{X}(t,a)[\nabla X(t,a)v]da$
$\int_{D}u(T,a)\cdot(X_{T})_{\sharp}(v\rho_{0})da-\int_{D}u_{0}(a)\cdot(v\rho_{0})da-\int_{0}^{T}\int_{D}\nabla p(t,a)(X_{\sharp}v)(a)da$
$\int_{D}u(T,a)\cdot(X_{T})_{\sharp}(v\rho_{0})da-\int_{D}u_{0}(a)\cdot(v\rho_{0})da+\int_{0}^{T}\int_{D}p(t,a)(\nabla\cdot v)\circ X^{-1}(a)da$
$\int_{D}u(T,a)\cdot(X_{T})_{\sharp}(v\rho_{0})da-\int_{D}u_{0}(a)\cdot(v\rho_{0})da$
where we integrate by part, use (12), change the variable and apply the following commutation relation between push-forward and divergence:
$\nabla\cdot(X_{\sharp}v)=(\nabla\cdot v)\circ X^{-1}.$
Hence, we conclude for all $v\in\tilde{\mathfrak{g}}:=T_{e}\mathscr{G}$, the quantity
$\mathscr{I}_{t}:=\int_{D}u(t,a)\cdot[X_{\sharp}(\rho_{0}v)](t,a)da$ (46)
is conserved. Typical example includes the case $d=2$ and $v=F(t,\rho_{0})\nabla^{\perp}\rho_{0}$ for any smooth function $F:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}$. Plugging such $v$ back to (46), we have
$\mathscr{I}_{t}=\int_{D}u(t,a)\nabla^{\perp}\rho(t,a)\rho(t,a)F(\rho(t,a))da,\quad\rho(t,a)=\rho_{0}(X_{t}^{-1}(a)).$
Assume that $\mathsf{Ran}(\rho_{t})=[0,c]$ (here $c$ is independent of $t$ since $\rho_{t}$ is a rearrangement of $\rho_{0}$), by coarea formula we have:
$\mathscr{I}_{t}=\int_{0}^{c}\alpha F(\alpha)\bigg(\int_{\{\rho_{t}=\alpha\}}u(t,a)\cdot\frac{\nabla^{\perp}\rho_{t}}{\lvert\nabla\rho_{t}\rvert}d\mathscr{H}^{1}\bigg)d\alpha=\int_{0}^{c}\alpha F(\alpha)\oint_{\{\rho_{t}=\alpha\}}u(t,a)\cdot d\ell d\alpha$
$\int_{0}^{c}\alpha F(\alpha)\Gamma(t,\alpha)d\alpha,\quad\Gamma(t,\alpha):=\oint_{\{\rho_{t}=\alpha\}}u(t,a)\cdot d\ell.$
Thanks to the conservation of $\mathscr{I}_{t}$ and arbitrariness of $F$, we conclude $\Gamma(t,\alpha)=\Gamma(0,\alpha)$ for all $t$ and $\alpha\in\mathsf{Ran}(\rho_{t})$.
Remark 3.4.
It is a natural question to ask that whether such vorticity formulation would lead the way to the study of finite time blow up of IIE in dimension $2$ and $3$. We remark that the Boussinesq system, which is a fluid equation with active tracers coupled via Buoyancy, is demonstrated in the recent breakthrough [19][20] to blow up in finite time in its local well-posedness class. However, in view of the vortex dynamics, the source term in IIE is much harder to control compared to the Boussinesq case. On the other hand, analysis of lifespan and continuation criteria in IIE based on the vorticity formulation is applicable, see the companion paper [35].
We now discuss the stream-vorticity formulation in torus. Assume $D=\mathbb{T}^{2}$, the key observation is that: For IIE, the conservation of linear momentum reads
$\frac{d}{dt}\int_{D}\rho udx=-\int_{D}u\cdot\nabla(\rho u)dx-\int_{D}\nabla pdx=0.$
Which follows from integrating the first term by part and using mean-free property of gradient fields on torus. However, mean of $u$ is not conserved.
By the above observation, WLOG we assume that $\rho_{0}u_{0}$ is mean zero. Abuse our notation a little bit temporarily, we let
$\eta=\rho u=\nabla^{\perp}\psi+\nabla q.$
Now, the dynamical equation for $\omega$ and $\rho$ keep unchanged:
$\partial_{t}\omega+u\cdot\nabla\omega+\{\frac{1}{2}\lvert u\rvert^{2},\rho\}=0,\quad\partial_{t}\rho+u\cdot\nabla\rho=0,$
where $\{\cdot,\cdot\}$ is the standard Poisson bracket in $2$-d.
The constitutive law becomes different: Taking curl on $\eta$ and making use of $\nabla\cdot u=0$, we have:
$\omega=\Delta\psi,\quad u=\rho^{-1}\nabla^{\perp}\psi+\rho^{-1}\nabla p\quad\Longrightarrow\quad u=\mathbf{P}_{\rho}(\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega).$
Hence, the stream-vorticity formulation of IIE on $\mathbb{T}^{2}$ reads:
$\left\{\begin{aligned} &\partial_{t}\omega+\mathscr{L}_{u}\omega+\{\frac{1}{2}\lvert u\rvert^{2},\rho\}=0;\\ &u=\mathbf{P}_{\rho}(\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega);\\ &\partial_{t}\rho+u\cdot\nabla\rho=0.\end{aligned}\right$. (47)
Supplemented by mean-zero condition of $\rho_{0}u_{0}$. Notice that the above constitutive law is a general law and reduces to the weighted Biot-Savart law in case that $D$ is a simply connected domain:
$\mathbf{P}_{\rho}(\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega)=\mathcal{K}_{\rho}\nabla\times(\rho\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega)=\mathcal{K}_{\rho}\omega=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\omega.$
3.3 Local Existence Via Vorticity Formulation
We recall that in the general dimension $d$, IIE is shown to be locally well-posed in $H^{s+1}\times H^{s+1}$ for $s>d/2$, and we refer the readers to [14][15] [13] for various well-posedness results in $L^{p}$-based Sobolev spaces and endpoint Besov spaces.
Here, we show the local existence of strong solutions of IIE in Sobolev space via the vorticity formulation (42). We choose to work on a smooth and simply connected domain $D\subset\mathbb{R}^{2}$ for simplicity, but the proof of local well-posedness on $\mathbb{T}^{2}$ using (47) is also available following the same line of argument. We also expect that the Lagrangian–Eulerian system (25) yields a unified proof of local well-posedness for strong solutions in all dimensions. For a similar Lagrangian approach to the local well-posedness of hydrodynamic models, we refer the interested reader to [12][10][33][39]. However, we will only focus on the proof through our vorticity formulation to highlight its utility.
We start with the following elementary elliptic lemma. The proof is standard and is therefore omitted.
Lemma 3.5.
Assume that $\omega\in H^{2}$ and $\rho\in H^{3}$ satisfies the assumption 2.1, then the solution $\psi$ of (43) satisfies the following estimate:
$\lVert\psi\rVert_{H^{4}}\lesssim P(\lVert\rho\rVert_{H^{3}})\lVert\omega\rVert_{H^{2}}.$ (48)
Here $P$ is some monotonically increasing polynomial function.
We now state our local existence result.
Theorem 3.6.
Assume $d=2$, Let $(\omega_{0},\rho_{0})\in H^{2}\times H^{3}$ and let $\rho$ satisfies the Assumption 2.1, then (42) admits a unique solution local in time.
Proof of Theorem 3.6.
We consider the following iterative scheme:
$\left\{\begin{aligned} &\partial_{t}\omega^{n+1}+{u^{n}}\cdot\nabla\omega^{n+1}+\nabla\frac{1}{2}\lvert u^{n}\rvert^{2}\times\nabla\rho^{n}=0\\ &\partial_{t}\rho^{n+1}+u^{n}\cdot\nabla\rho^{n+1}=0\\ &u^{n+1}=\mathcal{K}_{\rho}^{n+1}\omega^{n+1}:=\nabla^{\perp}\mathbf{L}_{\rho^{n+1}}^{-1}\omega^{n+1}\end{aligned}\right$. (49)
The proof is Lagrangian in nature, we will use Sobolev estimate of transport equation to show existence of the fixed point of system (49). The proof is divided into $3$ steps.
  • Step 1. We define
    $\mathcal{U}_{R}=\{(\omega,u,\rho)\in C_{t}H^{2}\times C_{t}H_{\sigma}^{3}\times C_{t}H^{3}:\lVert\omega\rVert_{C_{t}H_{x}^{2}}\leq R_{1},\lVert u\rVert_{C_{t}H_{x}^{3}}\leq R_{2},\lVert\rho\rVert_{C_{t}H_{x}^{3}}\leq R_{3}\}$
    We aim to show that for sufficiently small time $T$ and positive numbers $R_{1},R_{2},R_{3}$ that we will choose later, the sequence $\{(\omega^{n},u^{n},\rho^{n})\}$ generated by the iterative scheme (1) stays in $U_{R}$.
    First, notice that:
    $\rho^{n+1}(t,\cdot)=\rho_{0}\circ A_{t}^{n},\quad\text{where }\dot{X}_{t}^{n}=u_{t}^{n}(X_{t}^{n}),\quad A_{t}^{n}=(X_{t}^{n})^{-1}$
    and $1/\rho^{n+1}$ is also transported. Hence, we have
    $\lVert\rho_{t}^{n+1}\rVert_{H^{3}},\lVert 1/\rho_{t}^{n+1}\rVert_{H^{3}}\leq C(\rho_{0})\exp\bigg(C\int_{0}^{t}\lVert u^{n}\rVert_{H^{3}}d\tau\bigg)\leq C(\rho_{0})\exp(CtR_{2}).$
    Here, $C(\rho_{0})$ can be chosen as $C(\rho_{0})=\lVert\rho_{0}\rVert_{H^{3}}+\lVert\rho_{0}^{-1}\rVert_{H^{3}}$.
    Now we consider the estimate for vorticity equation. By Duhamel’s principle, we can write down the following Lagrangian solution formula:
    $\omega^{n+1}(t,\cdot)=\omega_{0}\circ A_{t}^{n}-\bigg[\int_{0}^{t}(\nabla^{*}u^{n}u^{n}\cdot\nabla^{\perp}\rho^{n})(X_{\tau}^{n})d\tau\bigg]\circ A_{t}^{n}$
    $\omega_{0}\circ A_{t}^{n}-\bigg[\nabla^{\perp}\rho_{0}\int_{0}^{t}\nabla^{*}X^{n}(\nabla^{*}u^{n}u^{n})(X_{\tau}^{n})d\tau\bigg]\circ A_{t}^{n}$
    Hence, we make use of the algebra structure of $H^{2}$ and standard result in propagation of Sobolev regularity along smooth flow (for instance, see [4][33]) to estimate:
    $\lVert\omega^{n+1}\rVert_{H^{2}}\leq\lVert\omega_{0}\circ A_{t}\rVert_{H^{2}}+\lVert\nabla\rho_{0}\circ A_{t}\rVert_{H^{2}}\exp(CtR_{2})\int_{0}^{t}\lVert\nabla X^{n}\rVert_{H^{2}}\lVert(\nabla^{*}uu)(X^{n})\rVert_{H^{2}}d\tau$
    $\leq\lVert\omega_{0}\rVert_{H^{2}}\exp(CtR_{2})+R_{3}\exp(2CtR_{2})\int_{0}^{t}\exp(2C\tau R_{2})R_{2}^{2}dt$
    $\leq\exp(CtR_{2})(\lVert\omega_{0}\rVert_{H^{2}}+CR_{2}^{2}R_{3}t)$ (50)
    Here we temporarily choose $T$ to be sufficiently small such that
    ${CR_{2}T}\leq 1,\quad CR_{2}^{2}R_{3}T\leq\lVert\omega_{0}\rVert_{H^{2}}$
    where $C>1$ is some generic large number. We leave $R_{2}$ to be fixed later. Thanks to this choice, for all $t\in[0,T]$ we have:
    $\lVert\rho^{n+1}\rVert_{H^{3}},\lVert 1/\rho^{n+1}\rVert_{H^{3}}\leq eC(\rho_{0}),\quad\lVert\omega_{t}^{n+1}\rVert_{H^{2}}\leq 2e\lVert\omega_{0}\rVert_{H^{2}}$
    Hence, we can choose $R_{1}=2e\lVert\omega_{0}\rVert_{H^{2}}$.
    Now we estimate $u^{n+1}$. Given $\lVert\omega^{n+1}\rVert\leq 2e\lVert\omega_{0}\rVert_{H^{2}}$ and $\lVert\rho^{n+1}\rVert_{H^{3}}\leq eC(\rho_{0})$, we have:
    $\lVert u^{n+1}\rVert_{H^{3}}\lesssim\lVert\mathbf{L}_{\rho^{n+1}}^{-1}\omega^{n+1}\rVert_{H^{4}}=\lVert\psi^{n+1}\rVert_{H^{4}}\leq 2eCP(eC(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}}$
    where $P$ is the same polynomial as in (48). Hence, we may choose:
    $R_{2}=2eCP(eC(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}},\quad R_{3}=eC(\rho_{0}).$
    And the above analysis demonstrates that our iteration scheme stays in $\mathcal{U}_{R}$.
  • Step 2. Now, we show that the above iterative sequence $\{(\omega^{n},u^{n},\rho^{n})\}\subset\mathcal{U}_{R}$ is Cauchy in $C_{t}H^{1}\times C_{t}H_{\sigma}^{2}\times C_{t}H^{2}$. We start by noticing for $u^{n}=\nabla^{\perp}\psi^{n}$:
    $\frac{\mathbf{L}_{\rho^{n+1}}\psi^{n+1}}{\rho^{n+1}}-\frac{\mathbf{L}_{\rho^{n}}\psi^{n}}{\rho^{n}}=\frac{\omega^{n+1}}{\rho^{n+1}}-\frac{\omega^{n}}{\rho^{n}}$
    Which implies:
    $\Delta(\psi^{n+1}-\psi^{n})+\nabla\log\rho^{n+1}\cdot\nabla(\psi^{n+1}-\psi^{n})+\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}=$
    $\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}+(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}.$
    Denote by $C_{R}$ constants depending on $R_{1},R_{2},R_{3}$ which may change from line to line. We now recall the following paraproduct inequality (see [38]): For $f\in H^{s}$, $g\in H^{r}$ with $r>s\geq 1$:
    $\lVert fg\rVert_{H^{s}}\leq\lVert f\rVert_{H^{s}}\lVert g\rVert_{H^{r}}.$ (51)
    By (51), we control the RHS as follows:
    $\bigg\lVert\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\omega^{n+1}-\omega^{n}\rVert_{H^{1}},\quad\bigg\lVert(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}},$
    and the commutator term satisfies
    $\lVert\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}}.$
    Hence, by standard elliptic estimate as in Lemma 3.5, let $\delta\omega^{n}:=\omega^{n+1}-\omega^{n}$, $\delta\rho^{n}:=\rho^{n+1}-\rho^{n}$ and $\delta u^{n}=u^{n+1}-u^{n}$, we conclude
    $\lVert\delta u^{n}\rVert_{H^{2}}\leq\lVert\psi^{n+1}-\psi^{n}\rVert_{H^{3}}\leq C_{R}(\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}).$ (52)
    Straightforward computation yields the following equation for $\delta\omega^{n}$ and $\delta\rho^{n}$:
    $\partial_{t}\delta\omega^{n}+u^{n}\cdot\nabla\delta\omega^{n}+\delta u^{n}\cdot\nabla\omega^{n}=\frac{1}{2}(\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\})=:\sigma_{n},$
    $\partial_{t}\delta\rho^{n}+u^{n}\cdot\nabla\delta\rho^{n}+\delta u^{n}\cdot\nabla\rho^{n}=0.$
    Again, using Duhamel’s formula we conclude:
    $\delta\omega^{n}=\delta\omega_{0}\circ A_{t}^{n}+\bigg(\int_{0}^{t}(\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n},$
    $\delta\rho^{n}=\delta\rho_{0}\circ A_{t}^{n}-\bigg(\int_{0}^{t}(\delta u^{n}\cdot\nabla\rho^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n}.$
    Follow the same line of computation as for $\omega^{n}$, since $\delta\omega_{0}=\delta\rho_{0}=0$, we have the following Sobolev estimates:
    $\lVert\delta\omega^{n}\rVert_{H^{1}}\leq e\int_{0}^{t}\lVert\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n}\rVert_{H^{1}}d\tau\leq e\int_{0}^{t}\lVert\sigma_{n}\rVert_{H^{1}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau;$
    $\lVert\delta\rho^{n}\rVert_{H^{2}}\leq e\int_{0}^{t}\lVert\delta u^{n}\cdot\nabla\rho^{n}\rVert_{H^{2}}d\tau\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau.$ (53)
    Now we estimate $\sigma_{n}$. Notice that
    $\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\}=\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n}\rvert^{2},\rho^{n-1}\}+\{\lvert u^{n}\rvert^{2},\rho^{n-1}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\}$
    $\{\lvert u^{n}\rvert^{2},\delta\rho^{n-1}\}+\nabla^{*}u^{n}(u^{n}-u^{n-1})\nabla^{\perp}\rho^{n-1}+\nabla^{*}(u^{n}-u^{n-1})u^{n-1}\nabla^{\perp}\rho^{n-1}$
    $\{\lvert u^{n}\rvert^{2},\delta\rho^{n-1}\}+\nabla^{*}u^{n}\delta u^{n-1}\nabla^{\perp}\rho^{n-1}+\nabla^{*}\delta u^{n-1}u^{n-1}\nabla^{\perp}\rho^{n-1}.$
    Hence, we conclude that
    $\lVert\sigma_{n}\rVert_{H^{1}}\leq C_{R}(\lVert\delta\omega^{n-1}\rVert_{H^{1}}+\lVert\delta\rho^{n-1}\rVert_{H^{2}})$
    and
    $\lVert\delta\omega^{n}\rVert_{H^{1}}\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n-1}\rVert_{H^{1}}+\lVert\delta\rho^{n-1}\rVert_{H^{2}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau.$ (54)
  • Step 3. Now we define: $E^{n}(t):=\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}$. By (52), (53), (54), we obtain:
    $E^{n}(t)\leq C_{R}\int_{0}^{t}E^{n}(\tau)d\tau+C_{R}\int_{0}^{t}E^{n-1}(\tau)d\tau,\quad E(0)=0$
    Now we denote by $U_{n}:=\sup_{t\in[0,T]}E^{n}(t)$ and rewrite the above inequality as
    $E^{n}(t)\leq C_{R}\int_{0}^{t}E^{n}(\tau)d\tau+C_{R}\int_{0}^{t}E^{n-1}(\tau)d\tau\leq C_{R}TU_{n}+C_{R}TU_{n-1}.$
    Therefore, taking supremum over $t$ on the LHS we conclude:
    $U_{n}\leq C_{R}TU_{n}+C_{R}TU_{n-1}.$
    Choosing $T$ smaller if necessary, such that $C_{R}T\leq 1/3$, we have
    $\frac{2}{3}U_{n}\leq(1-C_{R}T)U_{n}\leq C_{R}TU_{n-1}\leq\frac{1}{3}U_{n-1}.$
    Therefore, we obtain $U_{n}\leq U_{n-1}/2\leq 2^{-(n-1)}U_{1}$. In particular, we demonstrate the summability of the sequence
    $U_{n}=\sup_{t\in[0,T]}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}.$
    As a matter of fact, we conclude that $(\omega^{n},\rho^{n})$ is Cauchy in $C([0,T];H^{1}\times H^{2})$. Now, $u^{n}$ is also Cauchy in $C([0,T];H_{\sigma}^{2})$, thanks to (52). We therefore have a unique limit $(\omega,u,\rho)$.
    Finally, by our construction and thanks to Banach-Alaoglu theorem, the sequence $(\omega^{n},\rho^{n},u^{n})$ converges in weak-* sense in $\mathcal{U}_{R}$, whose limit must coincide with the strong $C([0,T];H^{1}\times H^{2}\times H^{2})$-limit $(\omega,u,\rho)\in\mathcal{U}_{R}$, which is the unique solution of (42).
4 Lagrangian Analyticity of Inhomogeneous Incompressible Fluids
Lagrangian trajectories of many incompressible hydrodynamic models are known to be locally analytic in time, as long as the initial data lies in the local well-posedness class. We refer the reader to [11][22][34][36][40] for detailed treatment on this subject. On the other hand, such a nice property fails to hold when we switch to the compressible models (see [23]), since the pressure is no longer an elliptic constraint, but forces the system to be hyperbolic conservation law with finite speed of propagation. Therefore, there may exist points which stay rest and then start to move, which prevents analyticity of the Lagrangian displacement. However, to the author’s knowledge, as an intermediate model between incompressible Euler and compressible Euler, whether such a property is enjoyed by IIE is not known in existing literature.
In this section, we approach the Lagrangian analyticity of the inhomogeneous incompressible Euler equation using our Lagrangian formulation. Our approach is similar to [40], where the vorticity formulation is used and Cauchy’s invariant plays a key role towards the iterative scheme for Taylor coefficients. In the present work, we stick to the momentum level, which allows us to deal with general dimensions in a unified way. The main theorem of this section is as follows.
Theorem 4.1.
Let $s>d/2$ and assume $u_{0}\in H_{\sigma}^{s+1}$, $\rho_{0}\in H^{s+1}$ satisfies assumption (A), then there exists a time threshold $T_{*}$ such that the Lagrangian flow $X$ of solution $u$ of IIE is analytic for $t\in(0,T_{*})$
It turns out that inhomogeneity doesn’t prevent analyticity, by comparing above theorem 4.1 with the compressible Euler counterexample in [23]. The crucial structure of incompressible fluids that leads to the analyticity is the hidden ellipticity in these models (The Biot-Savart law). Such a regularization mechanism, from the Lagrangian viewpoint, is manifested by the following pointwise integration by part lemma (cf. [12][10]).
Lemma 4.2.
Let $\ell,u\in H^{s+1}$, $s>d/2$. Then $\mathbf{P}(\nabla^{*}\ell u)\in H^{s+1}$.
Proof of Lemma 4.2.
It suffices to show $\partial_{i}\mathbf{P}(\nabla^{*}\ell u)\in H^{s}$ for all $i$. For each fixed $1\leq i\leq d$, by commutativity between Fourier multipliers we have
$\partial_{i}\mathbf{P}(\partial_{k}\ell_{j}u_{j})=\mathbf{P}(\partial_{ki}\ell u_{j}+\partial_{k}\ell_{j}\partial_{i}u_{j})=\mathbf{P}(\partial_{k}(\partial_{i}\ell_{j}u_{j})-\partial_{i}\ell_{j}\partial_{k}u_{j}+\partial_{k}\ell_{j}\partial_{i}u_{j})$
$\mathbf{P}(-\partial_{i}\ell_{j}\partial_{k}u_{j}+\partial_{k}\ell_{j}\partial_{i}u_{j})\in H^{s}$
where the last line follows from the fact that $H^{s}$ is an algebra and $\mathbf{P}$ is a 0-order multiplier. ∎
We now proceed to prove theorem 4.1.
Proof.
We will use our Lagrangian representation to obtain a recursive Taylor formula for the formal time series of Lagrangian displacement, and prove convergence of the formal time series. The proof is divided into $4$ steps.
  • Step 1.(Set up the Lagrangian Framework)
    Recall that in proof of theorem 3.1 (III) we derived:
    $\eta_{t}=\rho_{t}u_{t}+\nabla q_{t}=\nabla^{*}A_{t}\bigg(\eta_{0}-\int_{0}^{t}\frac{1}{2}\lvert\dot{X}_{\tau}(x)\rvert^{2}\nabla\rho_{0}(x)d\tau\bigg)\circ A_{t}$
    Acting $X_{t}$ to the above equality in the sense of pull-back of covector we have
    $\nabla^{*}X_{t}(\rho_{t}u_{t}+\nabla q_{t})\circ X_{t}=\eta_{0}-\nabla\rho_{0}\int_{0}^{t}\frac{1}{2}\lvert\dot{X}_{\tau}\rvert^{2}d\tau$
    Denote by $\ell_{t}(a):=X_{t}(a)-a$ the Lagrangian displacement, simplify the LHS, we end up with:
    $\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\nabla(q_{t}\circ X_{t})+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert(\tau)d\tau=\eta_{0},$ (55)
    which suggests that
    $\mathbf{P}\bigg[\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert^{2}(\tau)d\tau\bigg]=\mathbf{P}\eta_{0}.$ (56)
    Meanwhile, the following determinant identity relation automatically holds:
    $\det(\mathbf{Id}+\nabla\ell(t))=1\quad\text{for all }t\geq 0$ (57)
    Now, assuming for sufficiently small $t>0$, we have the following formal series expansion of $\ell$:
    $\ell_{j}(t)=\sum_{k=1}^{\infty}t^{k}\ell_{j}^{(k)},$
    here $\ell_{j}$ is the $j$-th component of $\ell$. Notice that here $\ell^{(0)}=0$, $\ell^{(1)}=u_{0}$. Straightforward computation yields:
    $(\nabla^{*}\ell(t))_{ij}=\sum_{k=1}^{\infty}t^{k}\partial_{i}\ell_{j}^{(k)},\quad\dot{\ell}_{j}(t)=\sum_{k=0}^{\infty}(k+1)t^{k}\ell_{j}^{(k+1)}.$
    Hence, we moreover have:
    $(\nabla^{*}\ell\dot{\ell})_{i}=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}(m-k+1)\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m+1-k)};$
    $\int_{0}^{t}\lvert\dot{\ell}(\tau)\rvert^{2}d\tau=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}\frac{k(m+1-k)}{m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}.$
    Combining the above together, we conclude that (56) reads for all $n\geq 1$:
    $\mathbf{P}(\rho_{0}u_{0})=$
    $\mathbf{P}\bigg[\rho_{0}\ell_{i}^{(1)}+\sum_{m=1}^{\infty}t^{m}\bigg((m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg)\bigg],$
    which implies that coefficient of $t^{m}$ vanishes at all order: For all $m\geq 1$
    $\mathbf{P}\bigg[(m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg]=0.$ (58)
    Moreover, the determinant identity (57) naturally implies
    $0=\log\det(\mathbf{Id}+\nabla\ell(t))=\mathsf{Tr}\log(\mathbf{Id}+\nabla\ell(t))=\mathsf{Tr}\bigg(\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}[\nabla\ell(t)]^{k}\bigg)$
    $\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\mathsf{Tr}\bigg[\bigg(\sum_{r=1}^{\infty}\nabla\ell^{(r)}t^{r}\bigg)^{k}\bigg].$
    Hence, the vanishing of coefficient for $t^{m}$ implies:
    $0=\sum_{k=1}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m,r_{i}\geq 1}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})}).$ (59)
  • Step 2. We now derive the iterative scheme for coefficients $\{\ell^{(k)}\}_{k\geq 1}$ based on (58)(59). To this end, we rewrite (58) as
    $(m+1)\ell_{i}^{(m+1)}\rho_{0}+U_{i}^{(m)}=\nabla p_{i}^{(m)}$ (60)
    where $U^{(m)}$ is a vector field quadratic in lower-order terms:
    $U^{(m)}=\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big).$
    Meanwhile, (59) implies:
    $\mathsf{Tr}(\nabla\ell^{(m)})=\nabla\cdot\ell^{(m)}=\sum_{k=2}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m,r_{i}\geq 1}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})})=:W^{(m)}$ (61)
    Here, the expansion eventually leads to a polynomial of lower order $\nabla\ell^{(r)}$s’ entries with degree at most $m$.
    Therefore, the following recursive relation holds: Fix $m$ and assume we already know $\{\ell^{(k)}\}_{k=1}^{m}$, then we can explicitly compute $U^{(m)}$, $W^{(m+1)}$ and $\nabla\cdot\ell^{(m+1)}=W^{(m+1)}$. Taking into account (60) we have:
    $\nabla\cdot\ell^{(m+1)}=\frac{1}{m+1}\nabla\cdot(\rho_{0}^{-1}(\nabla p^{(m)}-U^{(m)}))=W^{(m+1)},$
    which guarantees:
    $\nabla p^{(m)}=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}+\nabla\cdot(\rho_{0}^{-1}U^{(m)})].$ (62)
    Therefore one can produce $\ell^{(m+1)}$ from $\{\ell^{(k)}\}_{k=1}^{m}$ as follows:
    $\ell^{(m+1)}=\frac{1}{(m+1)\rho_{0}}\big[\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}U^{(m)})]-U^{(m)}\big]$
    Which is an explicit recursion formula for Taylor coefficients.
  • Step 3.(Elliptic Estimates) Introduce the following generating function:
    $\bar{\zeta}(t):=\sum_{k=1}^{\infty}t^{k}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}$ (63)
    We aim to show that the series (63) converges absolutely.
    Here, we consider the Leray-Hodge decomposition $U^{(m)}=\mathbf{P}U^{(m)}+\nabla\varphi^{(m)}$. Notice then
    $\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}U^{(m)})=\nabla\varphi^{(m)}+\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})$
    First, we need to show that $\ell^{(m+1)}\in H^{s+1}$, given $\{\ell^{(k)}\}_{k=1}^{m}\subset H^{s+1}$. Note that with our assumption 2.1, $\Delta_{\rho_{0}}$ is uniformly elliptic. The above equality (62) reads
    $\nabla(p^{(m)}-\varphi^{(m)})=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})]$ (64)
    and we have by (60):
    $\ell^{(m+1)}=\frac{1}{(m+1)\rho_{0}}\big[-\mathbf{P}U^{(m)}+\nabla(p^{(m)}-\varphi^{(m)})\big].$ (65)
    We first control $\mathbf{P}U^{(m)}$. We notice that:
    $\mathbf{P}U^{(m)}=\sum_{k=1}^{m}(m-1+k)\big(\mathbf{P}[\nabla^{*}\ell^{(k)}(\ell^{(m+1-k)}\rho_{0})]+\frac{k}{2m}\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})])$
    Therefore, $\mathbf{P}U^{(m)}\in H^{s+1}$ will hold true if we succeed in showing for any $1\leq k\leq m$:
    $\mathbf{P}[\nabla^{*}\ell^{(k)}(\ell^{(m+1-k)}\rho_{0})]\in H^{s+1},\quad\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]\in H^{s+1}.$
    Combining Lemma 4.2 and a simple manipulation on the second term:
    $\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]=-\mathbf{P}[\nabla(\ell^{(m+1-k)}\cdot\ell^{(k)})\rho_{0}]$
    $\mathbf{P}[\nabla^{*}\ell^{(m+1-k)}\cdot\ell^{(k)}\rho_{0}]-\mathbf{P}[\nabla^{*}\ell^{(k)}\cdot\ell^{(m+1-k)}\rho_{0}],$
    we eventually conclude the following estimate:
    $\lVert\mathbf{P}U^{(m)}\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}.$ (66)
    On the other hand, rewrite (62) as
    $\nabla(p^{(m)}-\varphi^{(m)})=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})].$
    Now, a slight variant of Lemma 3.5 (taking $\rho=1/\rho_{0}$) implies:
    $\lVert\nabla(p^{(m)}-\varphi^{(m)})\rVert_{H^{s+1}}\leq C(\rho_{0})(m+1)\lVert W^{(m+1)}\rVert_{H^{s}}+\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}.$
    Since $W^{(m+1)}$ consists of the sum of product of $\nabla\ell^{(r)}$ with $1\leq r\leq m$, the fact that $H^{s}$ is an algebra implies $W^{(m+1)}\in H^{s}$, with:
    $\lVert W^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1,r_{i}\geq 1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}},$ (67)
    while above estimate (66) implies
    $\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}.$
    Hence, we conclude from (65):
    $\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}+\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}},$ (68)
    where RHS is a polynomial of $\lVert\nabla\ell^{(r)}\rVert_{H^{s}}$ up to order $d$. Hence, there exists positive numbers $C_{2},\cdots,C_{d}$ such that
    $\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\leq\sum_{j=2}^{d}C_{j}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}.$ (69)
  • Step 4. (Conclude Analyticity) Now based on the recurrence relation, we can recursively build a sequence $\{\beta_{m}\}$ dominating coefficients $\{\lVert\nabla\ell^{(m)}\rVert_{H^{s}}\}$ of our generating function $\zeta$ as follows:
    $\beta_{1}=\lVert\nabla u_{0}\rVert_{H^{s}},\quad\beta_{m+1}=\sum_{j=2}^{d}C_{j}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\beta_{r_{i}}$
    and we define the formal series
    $\tilde{\zeta}(t)=\sum_{m=1}^{\infty}\beta_{m}t^{m}$
    which is the formal expansion of the unique root of the following polynomial:
    $P(t,\zeta)=-\zeta+t\lVert\nabla u_{0}\rVert_{H^{s}}+C_{2}\zeta^{2}+\cdots+C_{d}\zeta^{d}$ (70)
    To see the uniqueness of root of $P(t,\cdot)$, notice first that for each $t\geq 0$, solving $P(t,\zeta)=0$ is equivalent to solve
    $J_{t}(\alpha):=\alpha-\sum_{j=2}^{d}C_{j}\alpha^{j}=t\|\nabla u_{0}\rVert_{H^{s}}$
    We observe that for any $t>0$, $J_{t}(0)=0$, while $J_{t}^{\prime}(0)=1>0$ and $J_{t}(\alpha)\to$-$\infty$ as $\alpha\to\infty$. Hence, we consider the first $\alpha_{*}$ with sufficiently small derivative: Fix $0<\delta\ll 1$, we consider
    $\alpha^{*}=\min\{\alpha:J^{\prime}(\alpha)=\delta\},\quad J^{\prime}(\alpha)>\delta\text{ for all }0<\alpha<\alpha^{*}$
    We now define
    $\Psi_{t}(\zeta):=t\rVert\nabla u_{0}\rVert_{H^{s}}+\sum_{j=2}^{d}C_{j}\zeta^{j}$
    Notice that
    $\lvert\Psi_{t}^{\prime}(\zeta)\rvert\leq\sum_{j=2}^{d}jC_{j}\zeta^{j-1}=1-J^{\prime}(\zeta)\leq 1-\delta,\quad\text{for all }\zeta\in[0,\alpha_{*}]$
    Hence, for each $t$, $\Psi_{t}(\zeta)$ is a contraction. Moreover, we notice that $\Psi_{t}$ is increasing, therefore:
    $\Psi_{t}([0,\alpha])\subset[\Psi_{t}(0),\Psi_{t}(\alpha)]=\bigg[t\|\nabla u_{0}\rVert_{H^{s}},t\|\nabla u_{0}\rVert_{H^{s}}+\sum_{j=2}^{d}C_{j}\alpha^{j}\bigg]$
    To apply Banach fixed point theorem, it suffices to show:
    $t\|\nabla u_{0}\rVert_{H^{s}}+\sum_{j=2}^{d}C_{j}\alpha^{j}\leq\alpha$
    Hence it suffices to pick $t$ such that
    $t\leq T(\alpha):=\|\nabla u_{0}\rVert_{H^{s}}^{-1}J(\alpha)$
    Therefore, for all $t\in[0,T(\alpha)]$ with $\alpha\in(0,\alpha^{*})$, $\Psi_{t}$ has a unique fixed point. Hence, we find a unique $\zeta_{*}\in(0,\alpha^{*})$ such that $P(t,\zeta^{*})=0$.
    Now, since $P(t,\zeta)$ satisfies
    $P(0,0)=0,\quad\partial_{\zeta}P(0,0)=-1\neq 0$
    The implicit function theorem implies that locally around $(0,0)$, there exists an analytic function $\zeta_{*}(t)$ such that $P(t,\zeta_{*}(t))=0$. By uniqueness of root of $P(t,\cdot)$, we conclude the formal series solution must coincide with analytic function $\zeta_{*}(t)$, which concludes the absolute convergence of series $\tilde{\zeta}$ in $t\in(0,T(\alpha^{*}))$, hence also the convergence of $\bar{\zeta}(t)$. Hence, proof of the Lagrangian analyticity of IIE with convergence radius $t\in(0,T(\alpha^{*}))$ is complete.